### v3

# Small Gradient Norm Regret for Online Convex Optimization

Abstract
This paper introduces a new problem-dependent regret measure for online convex optimization with smooth losses. The notion, which we call the $G^{\star}$ regret, depends on the cumulative squared gradient norm evaluated at the decision in hindsight. We show that the $G^{\star}$ regret strictly refines the existing $L^{\star}$ (small loss) regret, and that it can be arbitrarily sharper when the losses have vanishing curvature around the hindsight decision. We establish upper and lower bounds on the $G^{\star}$ regret and extend our results to dynamic regret and bandit settings. As a byproduct, we refine the existing convergence analysis of stochastic optimization algorithms in the interpolation regime. Some experiments validate our theoretical findings.
1 Introduction
Online convex optimization (OCO) is a powerful framework for handling sequential decision-making problems [17, 25]. In an OCO problem, a learner makes sequential decisions over time horizon $T$ in an adversarial setting: at each time step $t=1,\ldots,T$, the learner chooses some decision $x^{t}$ from a convex set $\mathcal{X}\subseteq\mathbb{R}^{n}$ and receives a convex loss $\ell_{t}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ specified by the adversary. The goal of an OCO algorithm is to make decisions $\{x^{t}\}_{t=1}^{T}$ that achieve low cumulative regret with respect to an arbitrary decision $x\in\mathcal{X}$
$\rho_{T}(x)\coloneqq\sum_{t=1}^{T}\ell_{t}(x^{t})-\sum_{t=1}^{T}\ell_{t}(x),$
and in particular, the best decision in hindsight $x^{\star}\in\operatornamewithlimits{arg\,min}_{x\in\mathcal{X}}\sum_{t=1}^{T}\ell_{t}(x)$. Despite the powerful adversary, existing OCO algorithms can achieve sublinear regret under different assumptions on the losses $\ell_{t}$. For example, when the losses are convex and Lipschitz continuous, [41] shows that online gradient descent (OGD) achieves $\mathcal{O}(\sqrt{T})$ regret; if $\ell_{t}$ are additionally strongly convex or exponentially concave, then $\mathcal{O}(\log T)$ regret is attainable using OGD or online Newton step [16]. These regret bounds are minimax optimal in terms of the corresponding loss function classes [1].
Beyond the assumptions discussed above, another popular condition on the loss functions is smoothness, which is often satisfied by loss functions in machine learning tasks [29]. When the loss functions are non-negative and have Lipschitz continuous gradients, it is possible to obtain a problem-dependent regret $\mathcal{O}(\sqrt{L_{T}^{\star}})$ [29], where $L_{T}^{\star}=\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\ell_{t}(x)=\sum_{t=1}^{T}\ell_{t}(x^{\star})$ denotes the cumulative loss under the best decision in hindsight. This regret, often known as the $L^{\star}$-regret or small loss bound [25, 38, 34, 37], can be significantly sharper than the standard $\mathcal{O}(\sqrt{T})$ regret upper bound if $L_{T}^{\star}=\mathcal{O}(1)$, while still maintaining the worst-case $\mathcal{O}(\sqrt{T})$ regret. Recent work has developed parameter-free online learning algorithms that achieve $\mathcal{O}(\sqrt{L_{T}^{\star}})$ regret [38, 39] as well as guarantees in terms of other problem-dependent quantities.
Despite its strengths, the small loss regret ($L^{\star}$ regret) has certain limitations. First, most existing results for $L^{\star}$ regret assume that the losses are non-negative. Although this assumption can be removed by recentering each loss via $\bar{\ell}_{t}(x)\coloneqq\ell_{t}(x)-\inf_{x\in\mathbb{R}^{n}}\ell_{t}(x)$, it implicitly requires the losses to be lower bounded, a condition that is violated even by linear losses. Moreover, the lack of translation invariance (under the addition of a constant) of $L_{T}^{\star}$ remains conceptually unsatisfactory. These limitations motivate the search for an alternative regret measure that avoids these shortcomings.
Contributions.
This paper addresses the above problems by introducing the $G^{\star}$ regret, which depends on the quantity $G_{T}^{\star}\coloneqq\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{\star})\|^{2}$, the cumulative non-stationarity at the best decision in hindsight. In particular,
  • We show that the $G^{\star}$ regret is at least as tight as the $L^{\star}$ regret and overcomes all its drawbacks: it is translation invariant and does not require the losses to be lower bounded over $\mathbb{R}^{n}$. Moreover, we prove that the $G^{\star}$ regret can be arbitrarily sharper than the $L^{\star}$ regret when the losses have a small local smoothness constant around $x^{\star}$. We support these claims with numerical experiments.
  • We prove that existing algorithms that adapt to the $L^{\star}$ regret also adapt to the $G^{\star}$ regret, achieving an $\mathcal{O}(\sqrt{G_{T}^{\star}})$ bound with a matching lower bound. We further extend our analysis to dynamic regret and bandit settings, and showcase the implications of $G^{\star}$ regret in the context of stochastic optimization. We also show how to design smoothed versions of the loss functions to allow adaptation to the constrained small loss bound.
1.1 Related literature
We survey two existing problem-dependent regret bounds in OCO based on smoothness. For a comprehensive review of the problem-dependent regret bounds, see [39] and the references therein.
$L^{\star}$ regret.
The $L^{\star}$ regret, also known as the small loss bound, can provide a tighter regret upper bound when the decision in hindsight achieves a small cumulative loss
$L_{T}^{\star}\coloneqq\min_{x\in\mathcal{X}}\sum_{t=1}^{T}[\ell_{t}(x)-\inf_{u\in\mathbb{R}^{n}}\ell_{t}(u)]=\sum_{t=1}^{T}[\ell_{t}(x^{\star})-\inf_{u\in\mathbb{R}^{n}}\ell_{t}(u)],$ (1)
where $x^{\star}\in\operatornamewithlimits{arg\,min}_{x\in\mathcal{X}}\sum_{t=1}^{T}\ell_{t}(x)$. In the literature, it is often assumed that $\{\ell_{t}\}_{t=1}^{T}$ are non-negative, and an upper bound $\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\ell_{t}(x)$, is typically used. For ease of later comparison, this paper adopts the definition (1), the recentered version of the $L^{\star}$ regret. Existing OCO algorithms can achieve $\mathcal{O}(\sqrt{L^{\star}_{T}})$ regret for general smooth convex functions and $\mathcal{O}(\log L^{\star}_{T})$ regret when the loss is exponentially concave or strongly convex [34, 29]. Recently, the $L^{\star}$ regret has also been extended to the adaptive regret [36, 37].
Gradient variation bound.
Aside from the $L^{\star}$ regret, the following gradient variation bound [4, 23] is another widely adopted problem-dependent quantity. It often appears in the context of online learning with prediction, and depends on the maximum cumulative variation of the gradients over $\mathcal{X}$
$V_{T}\coloneqq\sum_{t=2}^{T}\max_{x\in\mathcal{X}}\|\nabla\ell_{t}(x)-\nabla\ell_{t-1}(x)\|^{2}.$ (2)
Existing OCO algorithms can achieve $\mathcal{O}(\sqrt{V_{T}})$ regret [4, 38, 39], and recent work has extended these results beyond the standard Lipschitz continuous gradient setting [35, 40].
1.2 Notations
Throughout the paper, we use $\|\cdot\|$ to denote the Euclidean norm and $\langle\cdot,\cdot\rangle$ to denote the Euclidean inner product. Given a closed convex set $\mathcal{X}$, $\Pi_{\mathcal{X}}[\cdot]$ denotes the orthogonal projection onto $\mathcal{X}$. The diameter of a set is defined by $\mathrm{diam}(\mathcal{X})\coloneqq\max_{x,y\in\mathcal{X}}\|x-y\|$. We use $\mathbb{S}_{n}\coloneqq\{x\in\mathbb{R}^{n}:\|x\|=1\}$ to denote the unit sphere centered at the origin. For brevity, we use $\inf\ell\coloneqq\inf_{x\in\mathbb{R}^{n}}\ell(x)$ to denote the infimum of $\ell$ over $\mathbb{R}^{n}$. We define $n]\coloneqq\{1,\ldots,n\}$.
2 Small gradient norm regret in online convex optimization
This section defines the $G^{\star}$ regret and compares it with the existing problem-dependent regret bounds. We also show that several existing OCO algorithms achieve $\mathcal{O}(\sqrt{G^{\star}_{T}})$ regret, accompanied by a matching lower bound. The following assumptions will be frequently invoked.
  1. A1:
    The losses $\ell_{t}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ are $L$-smooth and convex
    $\ell_{t}(x)-\ell_{t}(y)-\langle\nabla\ell_{t}(y),x-y\rangle\geq\tfrac{1}{2L}\|\nabla\ell_{t}(x)-\nabla\ell_{t}(y)\|^{2}\quad\text{for all }x,y\in\mathbb{R}^{n}.$ (3)
  2. A2:
    The closed convex set $\mathcal{X}$ is bounded with $\mathrm{diam}(\mathcal{X})\leq D$.
The assumption A1 looks slightly different from, but is equivalent to, the standard smoothness assumption $\ell_{t}(x)-\ell_{t}(y)-\langle\nabla\ell_{t}(y),x-y\rangle\leq\frac{L}{2}\|x-y\|^{2}$: an $L$-smooth convex function on $\mathbb{R}^{n}$ is $\frac{1}{L}$-strongly convex in the dual space. It yields a tighter lower bound than that provided by convexity alone [24]. This sharper lower bound is central to the derivation of the $G^{\star}$ regret.
Remark 1.
Deriving A1 from $L$-smoothness requires $\ell_{t}$ to be defined over the whole space. This assumption can be slightly relaxed to the case where $\mathcal{X}\subsetneq\text{int}\operatornamewithlimits{dom}\ell_{t}$ (see Section˜A.1). Note that the traditional setting of $L^{\star}$ regret also requires the losses to be defined over $\mathbb{R}^{n}$.
2.1 $G^{\star}$ regret bound
The definition of the $G^{\star}$ regret resembles that of the $L^{\star}$ regret, replacing the optimality gap of the loss $\ell_{t}(x)-\inf\ell_{t}$ by the non-stationarity $\|\nabla\ell_{t}(x)\|^{2}$:
Definition 2.1.
Given time horizon $T$ and a sequence of smooth convex losses $\{\ell_{t}\}_{t=1}^{T}$, the cumulative non-stationarity $G_{T}$ and the corresponding $G^{\star}$ regret are respectively defined as
$G_{T}(x)\coloneqq\sum_{t=1}^{T}\|\nabla\ell_{t}(x)\|^{2}\quad\text{and}\quad G^{\star}_{T}\coloneqq G_{T}(x^{\star})=\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{\star})\|^{2},$ (4)
where $x^{\star}\in\operatornamewithlimits{arg\,min}_{x\in\mathcal{X}}\sum_{t=1}^{T}\ell_{t}(x)$ is the minimizer of the cumulative losses.
The $G^{\star}$ regret inherits the translation invariance of the gradient and does not require the losses to be non-negative or lower bounded. In particular, for online linear optimization with $\ell_{t}(x)=\langle g_{t},x\rangle$,
$G_{T}^{\star}\equiv G_{T}(x)=\sum_{t=1}^{T}\|\nabla\ell_{t}(x)\|^{2}=\sum_{t=1}^{T}\|g_{t}\|^{2},$
and the $G^{\star}$ regret reduces to cumulative (sub)gradient norms, a quantity that many existing OCO algorithms adapt to. In contrast, the standard $L^{\star}$ regret is not well-defined in this setting.
Remark 2.
Although the OCO problem is constrained, the definition of the $G^{\star}$ regret does not involve the constraint set. It is consistent with the way that the $L^{\star}$ regret is defined: in the literature, a lower bound on $\ell_{t}$ over $\mathbb{R}^{n}$ (often 0) is used in place of $\ell_{t}$, and the $L^{\star}$ regret often comes in the shifted form $\sum_{t=1}^{T}[\ell_{t}(x)-\inf\ell_{t}]\leq\sum_{t=1}^{T}\ell_{t}(x)$.
Remark 3.
One may define the $G^{\star}$ regret via the gradient-norm minimization problem $\min_{x\in\mathcal{X}}G_{T}(x)$. However, the minimizer of $G_{T}$ need not coincide with the point $x^{\star}$ that minimizes the cumulative losses $\sum_{t=1}^{T}\ell_{t}$, and the resulting regret bounds would not be directly comparable to $L_{T}^{\star}$. For ease of comparison, we define $G_{T}^{\star}$ as the $x^{\star}$ that minimizes cumulative losses. This choice suffices to demonstrate the claimed tightness results.
Connection with existing regret.
The $G^{\star}$ regret measures the cumulative deviation from the unconstrained stationarity condition $\nabla\ell_{t}(x)=0$. It lower bounds the $L^{\star}$ regret (assuming it exists) by self-boundedness [29]:
$\|\nabla\ell_{t}(x)\|^{2}\leq 2L[\ell_{t}(x)-\inf\ell_{t}].$ (5)
Proposition˜2.1 relates the $G^{\star}$ regret to the existing problem-dependent regret bounds.
Proposition 2.1.
Under A1 and A2, $G^{\star}_{T}\leq 2L(L^{\star}_{T})$. Moreover, there exist loss sequences
  • $\{\ell_{t}^{A}\}_{t=1}^{T}$ such that $G^{\star}_{T}\ll L_{T}^{\star}$;
  • $\{\ell_{t}^{B}\}_{t=1}^{T}$ such that $V_{T}\ll L^{\star}_{T}~(\text{or }G^{\star}_{T})$;
  • $\{\ell_{t}^{C}\}_{t=1}^{T}$ such that $V_{T}\gg L^{\star}_{T}~(\text{or }G^{\star}_{T})$.
Proposition˜2.1 shows that small $L^{\star}$ regret implies small $G^{\star}$ regret, but the converse may not hold. Moreover, the $G^{\star}$ (or $L^{\star}$) regret and the gradient variation bound cannot imply each other. The rest of this paper focuses on the $L^{\star}$ and $G^{\star}$ regret. Below are examples where the $G^{\star}$ regret can be substantially smaller than $L^{\star}$. They all satisfy $\inf\ell_{t}=0$, and we drop $\inf\ell_{t}$ for brevity.
Example 2.1 (Logistic regression with cross-entropy loss).
Consider logistic regression with cross-entropy loss $\ell_{t}(x)=\log(1+\exp(-y_{t}\langle a_{t},x\rangle))$, where $\|a_{t}\|=1$ is a data vector and $y_{t}\in\{1,-1\}$. Then
$\nabla\ell_{t}(x)=\tfrac{-y_{t}a_{t}}{1+\exp(y_{t}\langle a_{t},x\rangle)}\quad\text{and}\quad\|\nabla\ell_{t}(x)\|^{2}\leq\ell_{t}(x)^{2}=o(\ell_{t}(x))$
whenever $\ell_{t}(x)\rightarrow 0$. The observation generalizes to multi-class classification.
Example 2.2 (Linear regression with $\ell_{p}$ loss).
Consider online linear regression with loss $\ell_{t}(x)=\tfrac{1}{p}|r_{t}(x)|^{p}$ for $p\geq 2$, where the residual $r_{t}(x)=\langle a_{t},x\rangle-b$. This loss is flat near the optimum, and the squared gradient norm vanishes faster than the function value: $\nabla\ell_{t}(x)=a_{t}|r_{t}(x)|^{p-2}r_{t}(x)$ and
$\tfrac{\|\nabla\ell_{t}(x)\|^{2}}{\ell_{t}(x)}=p\|a_{t}\|^{2}|r_{t}(x)|^{p-2}=p^{\frac{2p-2}{p}}\|a_{t}\|^{2}\ell_{t}(x)^{\frac{p-2}{p}}.$
Then $\|\nabla\ell_{t}(x)\|^{2}=o(\ell_{t}(x))$ whenever $r_{t}(x)\rightarrow 0$.
Example 2.3 (Exponential loss).
Consider exponential loss $\ell_{t}(x)=\mathrm{e}^{-\langle a_{t},x\rangle}$, where $\|a_{t}\|=1$ is a data vector. Then $\nabla\ell_{t}(x)=-\ell_{t}(x)a_{t}$ and $\|\nabla\ell_{t}(x)\|^{2}=\ell_{t}(x)^{2}=o(\ell_{t}(x))$ whenever $\ell_{t}(x)\rightarrow 0$.
Example 2.4 (Online regression with vanishing losses).
Consider online regression problem with loss $\ell_{t}(x)=\frac{1}{2}(a_{t}x)^{2},x\in[1,2]$. Suppose $a_{t}=t^{-p}$ for $p>0$. Then $\nabla\ell_{t}(x)=a_{t}^{2}x$ and $\|\nabla\ell_{t}(x)\|^{2}=t^{-4p}x^{2}=2t^{-2p}\ell_{t}(x)$. Hence $\|\nabla\ell_{t}(x)\|^{2}=o(\ell_{t}(x))$ as $t\rightarrow\infty$. In particular, when $p<\frac{1}{4}$, $L_{T}^{\star}=\frac{1}{2}\sum_{t=1}^{T}a_{t}^{2}=\Theta(T^{1-2p})$ and $G_{T}^{\star}=\Theta(T^{1-4p})$.
The above examples share the feature that the loss functions are flat in a neighborhood of the optimum. For such loss functions, the $G^{\star}$ regret is typically much tighter than the $L^{\star}$ regret in the (near-)interpolation regime where $x^{\star}$ is nearly optimal for each of the losses.
Remark 4.
When the losses satisfy a lower curvature bound such as $\mu$-strong convexity, the quantities $L_{T}^{\star}$ and $G_{T}^{\star}$ can be related via $2\mu L_{T}^{\star}\leq G_{T}^{\star}\leq 2LL_{T}^{\star}$. Hence, this paper focuses on the regime where $\mu$ is close to zero, for which the $\mathcal{O}(\tfrac{1}{\mu}\log T)$-type regret bounds become uninformative.
Remark 5.
The $L^{\star}$ regret can be obtained by directly assuming self-boundedness of the losses (5). Self-boundedness alone is weaker than $L$-smoothness and covers certain nonsmooth functions [25]. On the other hand, it is unclear whether (3) can be generalized to handle nonsmooth losses. We leave further investigations to future work.
2.2 Upper and lower bound for the $G^{\star}$ regret
This section presents algorithms that achieve an $\mathcal{O}(\sqrt{G^{\star}_{T}})$ regret upper bound with a matching lower bound. The key idea is simple. Consider the regret analysis of a prototypical online learning algorithm, such as online gradient descent (OGD) with constant learning rate $\eta>0$:
$x^{t+1}=\Pi_{\mathcal{X}}[x^{t}-\eta\nabla\ell_{t}(x^{t})].$
In the standard $L^{\star}$ regret analysis, we deduce, for any $x\in\mathcal{X}$, that
$\|x^{t+1}-x\|^{2}\leq{}$ $\|x^{t}-x\|^{2}-2\eta\langle\nabla\ell_{t}(x^{t}),x^{t}-x\rangle+\eta^{2}\|\nabla\ell_{t}(x^{t})\|^{2}$ (6)
$\leq{}$ $\|x^{t}-x\|^{2}-2\eta[\ell_{t}(x^{t})-\ell_{t}(x)]+\eta^{2}\|\nabla\ell_{t}(x^{t})\|^{2}$ (7)
$\leq{}$ $\|x^{t}-x\|^{2}-2\eta[\ell_{t}(x^{t})-\ell_{t}(x)]+2L\eta^{2}\cdot\ell_{t}(x^{t}),$ (8)
where (6) uses non-expansiveness of projection, (7) applies convexity $\ell_{t}(x)\geq\ell_{t}(x^{t})+\langle\nabla\ell_{t}(x^{t}),x-x^{t}\rangle$, and the $\|\nabla\ell_{t}(x^{t})\|^{2}$ term is further upperbounded in (8) using self-boundedness of (non-negative) $L$-smooth functions. However, a smooth function over $\mathbb{R}^{n}$ (A1) satisfies a stronger inequality
$\ell_{t}(x)\geq\ell_{t}(x^{t})+\langle\nabla\ell_{t}(x^{t}),x-x^{t}\rangle+\tfrac{1}{2L}\|\nabla\ell_{t}(x^{t})-\nabla\ell_{t}(x)\|^{2}.$ (9)
Using (9) allows us to prove a tighter bound compared to (7):
$\|x^{t+1}-x\|^{2}\leq\|x^{t}-x\|^{2}-2\eta[\ell_{t}(x^{t})-\ell_{t}(x)]\underbrace{-\tfrac{\eta}{L}\|\nabla\ell_{t}(x^{t})-\nabla\ell_{t}(x)\|^{2}+{\eta^{2}}\|\nabla\ell_{t}(x^{t})\|^{2}}_{\heartsuit}.\vskip-10.0pt$
Using the identity $\alpha\|a-b\|^{2}+\beta\|a\|^{2}=(\beta-\alpha)\big\|a-\tfrac{\alpha}{\alpha-\beta}b\big\|^{2}+\tfrac{\alpha\beta}{\alpha-\beta}\|b\|^{2},$ for $\eta\in(0,\frac{1}{L})$, we further bound $\heartsuit$ by:
$\heartsuit~={}$ $-\tfrac{\eta}{L}\|\nabla\ell_{t}(x^{t})-\nabla\ell_{t}(x)\|^{2}+\eta^{2}\|\nabla\ell_{t}(x^{t})\|^{2}$
$={}$ $(\eta^{2}-\tfrac{\eta}{L})\|\nabla\ell_{t}(x^{t})-\tfrac{1}{1-\eta L}\nabla\ell_{t}(x)\|^{2}+\tfrac{\eta^{2}}{1-\eta L}\|\nabla\ell_{t}(x)\|^{2}$
$\leq{}$ $\tfrac{\eta^{2}}{1-\eta L}\|\nabla\ell_{t}(x)\|^{2},$ (10)
where (10) uses $\eta^{2}-\tfrac{\eta}{L}\leq 0$ when $\eta\in(0,\frac{1}{L})$. We arrive at the following relation:
$\|x^{t+1}-x\|^{2}\leq\|x^{t}-x\|^{2}-2\eta[\ell_{t}(x^{t})-\ell_{t}(x)]+\tfrac{\eta^{2}}{1-\eta L}\|\nabla\ell_{t}(x)\|^{2}.$ (11)
A re-arrangement gives the standard regret inequality, and the regret bound follows by telescoping.
Theorem 2.1.
Under A1 and A2, OGD: $x^{t+1}=\Pi_{\mathcal{X}}[x^{t}-\eta\nabla\ell_{t}(x^{t})]$ with $\eta\in(0,\tfrac{1}{L})$ satisfies
$\rho_{T}(x)=\textstyle\sum_{t=1}^{T}\ell_{t}(x^{t})-\ell_{t}(x)\leq\tfrac{1}{2\eta}\|x^{1}-x\|^{2}+\tfrac{\eta}{2(1-\eta L)}\sum_{t=1}^{T}\|\nabla\ell_{t}(x)\|^{2}\leq\tfrac{D^{2}}{2\eta}+\tfrac{\eta}{2(1-\eta L)}G_{T}(x)$
for any $x\in\mathcal{X}$. In particular, choosing $x=x^{\star}$ and $\eta=\min\big\{\frac{D}{\sqrt{G^{\star}_{T}}},\tfrac{1}{2L}\big\}$ gives $\rho_{T}(x^{\star})\leq\max\{2LD^{2},\sqrt{2G^{\star}_{T}}D\}$.
The learning rate range in Theorem˜2.1 can be enlarged to $(0,\frac{2}{L})$ at the cost of re-introducing $L_{T}(x)$:
Corollary 2.1.
Under A1 and A2, OGD: $x^{t+1}=\Pi_{\mathcal{X}}[x^{t}-\eta\nabla\ell_{t}(x^{t})]$ with $\eta\in(0,\tfrac{2}{L})$ satisfies $\rho_{T}(x)\leq\tfrac{\|x^{1}-x\|^{2}}{(2-\eta L)\eta}+\tfrac{\eta}{2-\eta L}[LL_{T}(x)+\tfrac{1}{2-\eta L}G_{T}(x)]$ for any $x\in\mathcal{X}$.
Theorem˜2.1 resembles the guarantees of the $L^{\star}$ regret [25]. Using $G^{\star}_{T}\leq 2LL^{\star}_{T}$ from Proposition˜2.1, we obtain guarantees of the $L^{\star}$ regret: $\sum_{t=1}^{T}\ell_{t}(x^{t})-\ell_{t}(x)\leq\max\{2LD^{2},\sqrt{2LL^{\star}_{T}}D\}$. Besides, it always holds that $\sqrt{G^{\star}_{T}}=\mathcal{O}(\sqrt{T})$, retaining a worst-case $\mathcal{O}(\sqrt{T})$ regret.
Although Theorem˜2.1 achieves the desired regret guarantee, the learning rate $\eta$ must be chosen based on the unknown quantity $G^{\star}_{T}$, which depends on the future losses. This issue also arises for the $L^{\star}$ regret and can be addressed by employing adaptive online algorithms [32, 30]. These algorithms typically achieve an $\mathcal{O}(D\sqrt{\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t})\|^{2}})$ regret on the linearized losses, which, together with A1, yields
$\rho_{T}(x)={}\textstyle\sum_{t=1}^{T}\ell_{t}(x^{t})-\ell_{t}(x)\leq{}$ $\textstyle\sum_{t=1}^{T}\langle\nabla\ell_{t}(x^{t}),x^{t}-x\rangle-\frac{1}{2L}\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t})-\nabla\ell_{t}(x)\|^{2}$
$\leq{}$ $\textstyle\mathcal{O}(D\sqrt{\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t})\|^{2}})-\frac{1}{2L}\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t})-\nabla\ell_{t}(x)\|^{2}.$ (12)
Finally, Lemma˜2.1 below converts the regret upper bound of type (12) into the desired $\mathcal{O}(\sqrt{G_{T}(x)})$ regret guarantee by plugging in $a_{t}=\nabla\ell_{t}(x^{t})$ and $b_{t}=\nabla\ell_{t}(x)$.
Lemma 2.1.
Let $\{a_{t}\}_{t=1}^{T}$ and $\{b_{t}\}_{t=1}^{T}$ be two sequences of vectors and suppose $\alpha,\beta>0$. Then
$\alpha\sqrt{\sum_{t=1}^{T}\|a_{t}\|^{2}}-\beta\sum_{t=1}^{T}\|a_{t}-b_{t}\|^{2}\leq\alpha\sqrt{\sum_{t=1}^{T}\|b_{t}\|^{2}}+\tfrac{\alpha^{2}}{4\beta}.$
To showcase the analysis, we show that two adaptive algorithms from the literature achieve $\mathcal{O}(\sqrt{G_{T}(x)})$ regret.
AdaGrad-Norm.
The first one, AdaGrad-Norm [30], achieves the following regret guarantees.
Theorem 2.2.
Under A1 and A2, AdaGrad-Norm [30] (Algorithm˜2 in the appendix) $x^{t+1}=\Pi_{\mathcal{X}}[x^{t}-\eta_{t}\nabla\ell_{t}(x^{t})]$ with $\eta_{t}=\frac{\alpha}{\sqrt{\sum_{i=1}^{t}\|\nabla\ell_{i}(x^{i})\|^{2}}}$ and $\alpha=\tfrac{\sqrt{2}D}{2}$ satisfies $\rho_{T}(x)\leq\sqrt{2}\sqrt{G_{T}(x)}D+LD^{2}$ for any $x\in\mathcal{X}$.
AdaFTRL for linearized losses.
The second adaptive algorithm, AdaFTRL [26], achieves similar guarantees.
Theorem 2.3.
Under A1 and A2, suppose AdaFTRL [26] (Algorithm˜3 in the appendix) is applied to the linearized losses $\hat{\ell}_{t}(x)\coloneqq\langle\nabla\ell_{t}(x^{t}),x\rangle$. Further assume that AdaFTRL uses a $\lambda$-strongly convex regularizer $r$ with $\lambda\geq\frac{1}{2D^{2}}$ and $\max_{x\in\mathcal{X}}r(x)+1=:R$. Then it satisfies $\rho_{T}(x)\leq\sqrt{3}R\sqrt{G_{T}(x)}D+2R^{2}LD^{2}$ for any $x\in\mathcal{X}$.
Remark 6.
Since the conversion applies to any online algorithm with similar regret guarantees on the linearized losses, we can relax the bounded domain assumption using the recent results from unconstrained online learning (for example, [6]) and obtain a $\rho_{T}(x)\leq\mathcal{O}(\|x\|\sqrt{G_{T}(x)}+\max_{t\in[T]}\|\nabla\ell_{t}(x^{t})\|^{2}+(L+1)\|x\|^{2})$ regret result.
Lower bound.
Since $G^{\star}_{T}$ is equivalent to cumulative squared gradient norm for linear losses, the $\Omega(\sqrt{T})$ lower bound for online linear optimization implies an $\Omega(\sqrt{G^{\star}_{T}})$ regret lower bound.
Theorem 2.4.
Under A1 and A2, there exists a loss sequence $\{\ell_{t}\}_{t=1}^{T}$ such that for any online algorithm and $x\in\mathcal{X}$, we have $\rho_{T}(x)\geq\frac{1}{4}\sqrt{G_{T}^{\star}}D$.
3 Extensions
This section showcases extensions of the $G^{\star}$ regret to other OCO settings. Section˜3.1 covers an extension to the dynamic regret, and Section˜3.2 considers OCO with bandit feedback.
3.1 Dynamic $G^{\star}$ regret
Section˜2 focuses on the static regret with respect to a single competitor $x\in\mathcal{X}$. In this subsection, we consider dynamic regret, which allows for a sequence of competitors at the cost of a path length term in the final regret bound. Given a comparator sequence $\{\hat{x}_{t}\}_{t=1}^{T}$ with path length $\mathcal{P}_{T}(\{\hat{x}_{t}\}_{t=1}^{T})\coloneqq\sum_{t=2}^{T}\|\hat{x}_{t}-\hat{x}_{t-1}\|$, the dynamic regret with respect to this sequence is defined as
$\hat{\rho}(\{\hat{x}_{t}\}_{t=1}^{T})\coloneqq\sum_{t=1}^{T}\ell_{t}(x^{t})-\sum_{t=1}^{T}\ell_{t}(\hat{x}_{t}).$
As in the static regret setting, problem-dependent dynamic regret bounds have been established for both the $L^{\star}$ regret and the gradient variation bound [39, 38]. In particular, the dynamic version of the $L^{\star}$ regret is given by $\hat{L}_{T}(\{\hat{x}_{t}\}_{t=1}^{T})\coloneqq\sum_{t=1}^{T}[\ell_{t}(\hat{x}_{t})-\inf\ell_{t}]$. Online gradient descent with constant learning rate $\eta$ achieves an $\mathcal{O}(\tfrac{1+\mathcal{P}_{T}}{\eta}+\eta\hat{L}_{T})$ regret bound [38]. We now consider a dynamic analog of the $G^{\star}$ regret:
$\hat{G}_{T}(\{\hat{x}_{t}\}_{t=1}^{T})\coloneqq\sum_{t=1}^{T}\|\nabla\ell_{t}(\hat{x}_{t})\|^{2}.$
As in the static regret case, self-boundedness implies $\hat{G}_{T}(\{\hat{x}_{t}\}_{t=1}^{T})\leq 2L\hat{L}_{T}(\{\hat{x}_{t}\}_{t=1}^{T})$ given the same competitor sequence. Theorem˜3.1 below establishes a dynamic regret bound in $\hat{G}_{T}$.
Theorem 3.1.
Under A1 and A2, online gradient descent with $\eta\in(0,\tfrac{1}{4L}]$ satisfies $\hat{\rho}(\{\hat{x}_{t}\}_{t=1}^{T})\leq\tfrac{D(D+2\mathcal{P}_{T})}{2\eta}+\eta\hat{G}_{T}$ for any comparator sequence $\{\hat{x}_{t}\}_{t=1}^{T},~\hat{x}_{t}\in\mathcal{X}$. In particular, taking $\eta=\min\{\sqrt{\tfrac{{D(D+2\mathcal{P}_{T}})}{{2\hat{G}_{T}}}},\frac{1}{4L}\}$ gives $\hat{\rho}(\{\hat{x}_{t}\}_{t=1}^{T})\leq\max\{2LD(D+2\mathcal{P}_{T}),\sqrt{{2D(D+2\mathcal{P}_{T})\hat{G}_{T}}}\}=\mathcal{O}(\sqrt{(1+\mathcal{P}_{T}+\hat{G}_{T})(1+\mathcal{P}_{T})})$.
Again, Theorem˜3.1 is restrictive since the learning rate $\eta$ has to be chosen based on unknown quantities $\mathcal{P}_{T}$ and $\hat{G}_{T}$. This issue can be addressed by adopting the meta-expert algorithm framework in [39].
Theorem 3.2.
Under the same conditions as Theorem˜3.1, there exists an online algorithm that achieves $\mathcal{O}(\sqrt{(1+\mathcal{P}_{T}+\hat{G}_{T})(1+\mathcal{P}_{T})}\log\log T)$ dynamic regret without knowing $\hat{G}_{T}$ or $\mathcal{P}_{T}$.
The algorithm description and the proof of Theorem˜3.2 are left to the appendix.
3.2 $G^{\star}$ regret under bandit feedback
This subsection considers the setting in which access to the loss sequence $\{\ell_{t}\}_{t=1}^{T}$ is restricted to the two-point function-value feedback model [10, 9, 27, 18]. In this model, at each time step $t$, we draw a random direction $s_{t}$ uniformly from the unit sphere $\mathbb{S}_{n}$ and query the loss at two points: $(y^{t}_{+},y^{t}_{-})=(x^{t}+\mu s_{t},x^{t}-\mu s_{t})$, where $\mu>0$ is called the smoothing parameter. After observing $\ell_{t}(y^{t}_{+})$ and $\ell_{t}(y^{t}_{-})$, we construct the gradient estimator
$g_{t}\coloneqq\frac{n}{2\mu}\bigl[\ell_{t}(y^{t}_{+})-\ell_{t}(y^{t}_{-})\bigr]s_{t},$ (13)
and perform bandit gradient descent (BGD): $x^{t+1}=\Pi_{\mathcal{X}}[x^{t}-\eta_{t}g_{t}]$. Here, $g_{t}$ is not an unbiased estimator of the true gradient $\nabla\ell_{t}(x^{t})$; instead, it is an unbiased estimator of the gradient of $\tilde{\ell}_{t}$, a smoothed version of $\ell_{t}$, at $x^{t}$:
$\mathbb{E}_{s_{t}\sim\mathbb{S}_{n}}[g_{t}|x^{t}]=\nabla\tilde{\ell}_{t}(x^{t}),\quad\text{where}\quad\tilde{\ell}_{t}(x)\coloneqq\mathbb{E}_{s_{t}\sim\mathbb{S}_{n}}[\ell_{t}(x+\mu s_{t})].$ (14)
This smoothed function $\tilde{\ell}_{t}$ naturally serves as a surrogate function when performing BGD. Key properties of the smoothed function used in our analysis are summarized in the appendix. The following theorems establish regret bounds for BGD under both constant and adaptive learning rates.
Theorem 3.3.
Under A1, A2 and suppose $n\geq 8$, bandit gradient descent (Algorithm˜5 in the appendix) $x^{t+1}=\Pi_{\mathcal{X}}[x^{t}-\eta g_{t}]$ with constant learning rate $\eta\in(0,\tfrac{1}{4nL})$ satisfies
$\mathbb{E}[\rho_{T}(x)]\leq\textstyle\tfrac{D^{2}}{2\eta}+\tfrac{4n\eta}{1-4n\eta L}G_{T}(x)+\tfrac{4n\eta}{1-4n\eta L}TL^{2}\mu^{2}+\tfrac{\eta T}{2}n^{2}L^{2}\mu^{2}+\tfrac{TL\mu^{2}}{2}$
for any $x\in\mathcal{X}$. In particular, choosing $x=x^{\star}$, $\eta=\min\big\{\tfrac{D}{4\sqrt{nG^{\star}_{T}}},\tfrac{1}{8nL}\big\}$ and $\mu=\tfrac{D}{\sqrt{2nT}L}$ gives $\mathbb{E}[\rho_{T}(x^{\star})]\leq\max\{8nLD^{2},4D\sqrt{nG^{\star}_{T}}\}+\tfrac{LD^{2}}{2}$.
Theorem 3.4.
Under the same assumptions as Theorem˜3.3, bandit gradient descent (Algorithm˜6 in the appendix) $x^{t+1}=\Pi_{\mathcal{X}}[x^{t}-\eta_{t}g_{t}]$ with AdaGrad-Norm learning rate $\eta_{t}=\tfrac{\alpha}{\sqrt{\sum_{i=1}^{t}\|g_{i}\|^{2}}}$, $\alpha=\tfrac{\sqrt{2}D}{2}$ and $\mu=\tfrac{D}{n\sqrt{T}}$ satisfies $\mathbb{E}[\rho_{T}(x)]\leq\max\{16LD^{2}n,4D\sqrt{nG_{T}(x)}\}+5LD^{2}$ for any $x\in\mathcal{X}$.
4 Applications
We conclude by discussing the implications of the $G^{\star}$ regret.
4.1 Stochastic optimization under interpolation
Consider the following optimization problem
$\min_{x\in\mathcal{X}}~f(x)\coloneqq\mathbb{E}_{\xi\sim\Xi}[f(x,\xi)],$ (15)
where we assume that $f(x,\xi)$ is convex and $L$-smooth for all $\xi\sim\Xi$. The convergence analysis of stochastic gradient descent relies on assumptions on the stochastic noise, and the following two quantities, known as function noise and gradient noise respectively [13, 14], are widely used in the literature.
$(\sigma_{f}^{\star})^{2}\coloneqq\inf_{x}\mathbb{E}[f(x,\xi)-\inf_{u}f(u,\xi)]\quad\text{and}\quad(\sigma_{g}^{\star})^{2}\coloneqq{\mathbb{E}[\|\nabla f(x^{\star},\xi)\|^{2}]},$
where $x^{\star}\in\operatornamewithlimits{arg\,min}_{x}f(x)$ is assumed to exist. We make the following simplifying assumption.
  1. A3:
    There is some unconstrained minimizer $x^{\star}\in\operatornamewithlimits{arg\,min}_{x}f(x)$ such that $x^{\star}\in\mathcal{X}$.
Under A3, we have $(\sigma_{f}^{\star})^{2}=\min_{x\in\mathcal{X}}\mathbb{E}[f(x,\xi)-\inf_{u}f(u,\xi)]$, and it is easy to verify that $\sigma_{f}^{\star}$ and $\sigma_{g}^{\star}$ mirror the $L^{\star}$ and $G^{\star}$ regret, respectively. Let $\{\xi^{t}\}_{t=1}^{T}$ be i.i.d. samples from distribution $\Xi$. Taking $\ell_{t}(x)=f(x,\xi^{t})$ gives
$\mathbb{E}[L_{T}^{\star}]={}$ $\textstyle\mathbb{E}[\sum_{t=1}^{T}\ell_{t}(x^{\star})-\inf\ell_{t}]=\mathbb{E}[\sum_{t=1}^{T}f(x^{\star},\xi^{t})-\inf f(u,\xi^{t})]=T(\sigma_{f}^{\star})^{2},$
$\mathbb{E}[G_{T}^{\star}]={}$ $\textstyle\mathbb{E}[\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{\star})\|^{2}]=\mathbb{E}[\sum_{t=1}^{T}\|\nabla f(x^{\star},\xi^{t})\|^{2}]=T(\sigma_{g}^{\star})^{2}.$
When $\sigma_{f}^{\star}$ or $\sigma_{g}^{\star}$ is close to 0, the regime is known as (near-)interpolation [22]. Interpolation is a central setting in optimization for machine learning and has recently attracted significant attention [29, 21, 33, 2]. Since online algorithms with sublinear regret can be converted into an optimization algorithm via online-to-batch conversion [25], our $G^{\star}$ regret bounds from Section˜2.2 imply convergence guarantees for stochastic gradient methods and stochastic dual averaging (FTRL for linearized losses).
Corollary 4.1.
Suppose $f(x,\xi)$ is $L$-smooth and convex for all $\xi$ and that A2, A3 hold. Let $\ell_{t}(x)=f(x,\xi^{t})$ and $\{\xi^{t}\}_{t=1}^{T}$ be i.i.d. samples drawn from $\Xi$. Then
  • if OGD in Theorem˜2.1 is applied to $\{\ell_{t}\}_{t=1}^{T}$, then for any $\eta\in(0,\tfrac{1}{L})$, we have
    $\mathbb{E}[f(\bar{x}^{T})-f(x^{\star})]\leq\tfrac{1}{2\eta T}\|x^{1}-x^{\star}\|^{2}+\tfrac{\eta}{2(1-\eta L)}(\sigma_{g}^{\star})^{2};$
  • if AdaGrad-Norm in Theorem˜2.2 is applied to $\{\ell_{t}\}_{t=1}^{T}$, then $\mathbb{E}[f(\bar{x}^{T})-f(x^{\star})]\leq\tfrac{LD^{2}}{T}+\tfrac{\sqrt{2}D}{\sqrt{T}}\sigma_{g}^{\star}$;
  • if AdaFTRL from Theorem˜2.3 is applied to $\{\ell_{t}\}_{t=1}^{T}$, then $\mathbb{E}[f(\bar{x}^{T})-f(x^{\star})]\leq\tfrac{2R^{2}LD^{2}}{T}+\tfrac{\sqrt{3}RD}{\sqrt{T}}\sigma_{g}^{\star}$.
where $\bar{x}^{T}\coloneqq\tfrac{1}{T}\sum_{t=1}^{T}x^{t}$ is the average iterate and $R$ is defined in Theorem˜2.3.
Corollary˜4.1 improves the existing results [29] by replacing $\sigma_{f}^{\star}$ by $\sigma_{g}^{\star}$. Compared with the recent last-iterate convergence results of stochastic gradient descent [2, 12], our result enables adaptivity to smoothness and noise for the average iterate (Table˜1). Furthermore, using the anytime online-to-batch conversion from [5], similar last-iterate convergence follows immediately.
Corollary 4.2.
Under the same assumptions as Corollary˜4.1, algorithms in Corollary˜4.1 with anytime online-to-batch conversion [5] give the same guarantees for the last iterate $x^{T}$.
Table 1: Implications of the $G^{\star}$ regret for stochastic smooth convex optimization. Prefix AT denotes algorithm variants with anytime online-to-batch conversion [5]. The settings with bounded domain also assume that A3.
Reference Setting Algorithm Adapt to $\sigma^{\star}$ Rate
[29] Avg. iter, $\mathcal{X}=\mathbb{R}^{n}$ SGD No $\mathcal{O}(\frac{1}{T}+\frac{\sqrt{L}}{\sqrt{T}}\sigma_{f}^{\star})$
[2, 12] Last iter, $\mathcal{X}=\mathbb{R}^{n}$ SGD No $\mathcal{O}\big(\frac{\log T}{T}+\sqrt{\frac{\log T}{T}}\,\sigma_{g}^{\star}\big)$
[32] Avg. iter, $\mathrm{diam}(\mathcal{X})\leq D$ AdaGrad Yes $\mathcal{O}(\frac{1}{T}+\frac{\sqrt{L}}{\sqrt{T}}\sigma_{f}^{\star})$
This paper Avg. iter, $\mathrm{diam}(\mathcal{X})\leq D$ AdaGrad-Norm Yes $\mathcal{O}(\frac{1}{T}+\frac{1}{\sqrt{T}}\,\sigma_{g}^{\star})$
AdaFTRL
Last iter, $\mathrm{diam}(\mathcal{X})\leq D$ AT-AdaGrad-Norm Yes $\mathcal{O}(\frac{1}{T}+\frac{1}{\sqrt{T}}\,\sigma_{g}^{\star})$
AT-AdaFTRL
Improved analysis of AdaGrad-Norm.
As a byproduct, we also slightly refine the existing analysis of AdaGrad-Norm for unconstrained smooth convex optimization [20].
Theorem 4.1.
Suppose AdaGrad-Norm is applied to (15) with $\mathcal{X}=\mathbb{R}^{n}$ and that there exists $x^{\star}$ such that $\nabla f(x^{\star},\xi)=0$ surely. Further assume that each $f(x,\xi)$ is $L$-smooth. Then $\mathbb{E}[f(\bar{x}^{T})-f(x^{\star})]\leq\tfrac{L}{2T}(\tfrac{\|x^{1}-x^{\star}\|^{2}}{2\alpha}+\alpha\log(\tfrac{\sqrt{\mathrm{e}}\alpha L}{\|\nabla f(x^{1},\xi^{1})\|})+\tfrac{1}{2L}\|\nabla f(x^{1},\xi^{1})\|)^{2}$, where $\bar{x}^{T}\coloneqq\tfrac{1}{T}\sum_{t=1}^{T}x^{t}$ is the average iterate.
4.2 (Towards) constraint-aware $L^{\star}$ (and $G^{\star}$) regret via smooth surrogate loss
Although this paper focuses on constrained OCO, the definition of $L^{\star}_{T}$ measures suboptimality over $\mathbb{R}^{n}$ and does not take into account the constraint set $\mathcal{X}$. A more desirable measure is
$L^{\mathcal{X}}_{T}(x)\coloneqq\sum_{t=1}^{T}\ell_{t}(x)-\min_{u\in\mathcal{X}}\ell_{t}(u),$
which measures suboptimality relative to the constrained optimum rather than the unconstrained optimum $\sum_{t=1}^{T}\ell_{t}(x)-\inf\ell_{t}$. Similarly for $G^{\star}$, the gradient norm $\|\nabla\ell_{t}(x)\|$ is oblivious of the constraint set $\mathcal{X}$. Instead, a more appropriate measure should involve $\mathrm{dist}(-\nabla\ell_{t}(x),\mathcal{N}_{\mathcal{X}}(x))$, where $\mathcal{N}_{\mathcal{X}}(x)\coloneqq\{g:\langle g,x-y\rangle\geq 0,\text{ for all }y\in\mathcal{X}\}$ denotes the normal cone of $\mathcal{X}$ at $x$. The constrained cumulative suboptimality $L^{\mathcal{X}}_{T}$ can provide sharper problem-dependent constants. For example, suppose the unconstrained optimum is not in $\mathcal{X}$ but the constrained optimum $x^{\star}\in\operatorname{bd}\mathcal{X}$ is also the constrained optimum for most losses individually: $\ell_{t}(x^{\star})=\min_{x\in\mathcal{X}}\ell_{t}(x)$. In this case, $L^{\mathcal{X}}_{T}(x^{\star})=\mathcal{O}(1)$, whereas $L^{\star}_{T}$ can still be $\Omega(T)$. The rest of this section demonstrates the (im)possibility of generalizing the existing small loss (gradient) bound to adapt to the constraint set $\mathcal{X}$.
Hardness.
As discussed, the naive goal is to design an online algorithm that directly adapts to $L^{\mathcal{X}}_{T}$ (or $G^{\mathcal{X}}_{T}$) by replacing the traditional $\mathcal{O}(LD^{2}+D\sqrt{LL^{\star}_{T}})$ by $\mathcal{O}(LD^{2}+D\sqrt{LL^{\mathcal{X}}_{T}(x^{\star})})$. However, Proposition˜4.1 shows that this goal is generally not achievable.
Proposition 4.1.
For any $\tau\in[0,\tfrac{1}{6})$, there exists a loss sequence $\{\ell_{t}\}_{t=1}^{T}$ satisfying A1, A2 and that $\inf\ell_{t}=0$ such that for any online algorithm and $x\in\mathcal{X}$, $\frac{\rho_{T}(x)}{LD^{2}+D\sqrt{LL^{\mathcal{X}}_{T}(x)}}=\Omega(T^{\tau})$.
Despite the hardness result, whenever it is possible to redefine the losses, we show a strategy that enables dependence on $L^{\mathcal{X}}_{T}$ and $G^{\mathcal{X}}_{T}$ by replacing the losses by their smooth surrogates.
Easiness with smoothing.
Given a lower-semicontinuous convex function $\ell:\mathcal{X}\rightarrow\mathbb{R}$, its associated Moreau envelope with parameter $\gamma>0$, denoted by $\ell^{1/\gamma}(x)$, is given by
$\ell^{1/\gamma}(x)\coloneqq\min_{y\in\mathcal{X}}~\{\ell(y)+\tfrac{\gamma}{2}\|y-x\|^{2}\}.$
Regardless of the domain or smoothness of $\ell$, the function $\ell^{1/\gamma}$ is always $\gamma$-smooth and convex, with gradient given by $\nabla\ell^{1/\gamma}(x)=\gamma(x-\operatorname{prox}_{\ell/\gamma}(x))$, where $\operatorname{prox}_{\ell/\gamma}(x)=\operatornamewithlimits{arg\,min}_{x}\{\ell(y)+\tfrac{\gamma}{2}\|y-x\|^{2}\}$ is the proximal point. In particular, $\ell^{1/\gamma}$ and $\ell$ share the same minimizers and stationarity points, and $\nabla\ell^{1/\gamma}(x)$ approximates the quantity $\mathrm{dist}(-\nabla\ell(x),\mathcal{N}_{\mathcal{X}}(x))$ [7]. Hence, we similarly define $G^{\mathcal{X}}_{T}(x)\coloneqq\sum_{t=1}^{T}\|\nabla\ell_{t}^{1/\gamma}(x)\|^{2}$. In view of the properties of the smooth approximation, instead of solving the original OCO problem, suppose we consider the OCO problem defined on the smooth surrogates:
$\rho_{T}^{1/\gamma}(x)\coloneqq\sum_{t=1}^{T}\ell^{1/\gamma}_{t}(x^{t})-\ell^{1/\gamma}_{t}(x),\quad x^{t}\in\mathcal{X},t\in[T].$
Since each loss $\ell^{1/\gamma}_{t}$ is convex and $\gamma$-smooth, the results from the previous sections directly apply.
Proposition 4.2.
We have $G^{\mathcal{X}}_{T}(x)\leq 2\gamma L^{\mathcal{X}}_{T}(x)$, and the results from previous sections hold for $\rho^{1/\gamma}_{T}$. In particular, there exist online algorithms satisfying $\rho_{T}^{1/\gamma}(x)=\mathcal{O}(\gamma D^{2}+\sqrt{G^{\mathcal{X}}_{T}(x)})=\mathcal{O}(\gamma D^{2}+\sqrt{\gamma L^{\mathcal{X}}_{T}(x)})$ for any $x\in\mathcal{X}$.
We make two remarks on running online learning algorithms on the surrogate loss sequence $\{\ell^{1/\gamma}_{t}\}_{t=1}^{T}$.
First and importantly, the regret guarantee in Proposition˜4.2 is with respect to $\rho_{T}^{1/\gamma}$ instead of $\rho_{T}$, and the gap between $\ell^{1/\gamma}(x^{t})$ and $\ell(x^{t})$ can be up to $\mathcal{O}(\gamma^{-1})$ [3, 7]. In particular, using Proposition 4.2 of [7], we have
$\ell^{1/\gamma}(x)+\tfrac{1}{2\gamma}\|\nabla\ell^{1/\gamma}(x)\|^{2}\leq\ell(x)\leq\ell^{1/\gamma}(x)+\tfrac{1}{\gamma}\|\nabla\ell(x)\|^{2},$ (16)
and combining (16) with Proposition˜4.2 yields a bound on the regret:
$\rho_{T}(x)\leq\rho^{1/\gamma}_{T}(x)+\frac{1}{\gamma}\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t})\|^{2}\leq\mathcal{O}\big(\gamma D^{2}+\frac{1}{\gamma}\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t})\|^{2}+\sqrt{G_{T}^{\mathcal{X}}(x)}\big).$ (17)
Assuming $\|\nabla\ell(x^{t})\|\leq M$, the first two terms $\gamma D^{2}+\frac{1}{\gamma}TM^{2}$ require $\gamma=\Omega(\sqrt{T})$ to guarantee $\rho_{T}(x)\leq\mathcal{O}(\sqrt{T})$, and there is no regret improvement in the worst-case. In the appendix, we show the hardness of obtaining an $\mathcal{O}(LD^{2}+D\sqrt{G^{\mathcal{X}}_{T}(x)})$ on $\rho_{T}(x)$ using an argument similar to Proposition˜4.1. While this argument contains an improved bound for the smoothed loss, it does not always yield an improved bound for the original loss.
Remark 7.
The $G^{\star}$ regret can be strictly tighter than the $L^{\star}$ regret. To see this, notice that for $\gamma=\Theta(\sqrt{T})$, the $\mathcal{O}(\gamma D^{2}+\sqrt{\gamma L^{\mathcal{X}}_{T}(x)})$ upper bound yields a $\sqrt{\gamma L^{\mathcal{X}}_{T}(x)}$ term that can be as large as $\Theta(T^{3/4})$ in the worst case.
Second, running OGD on $\ell^{1/\gamma}_{t}$ requires computing $\nabla\ell_{t}^{1/\gamma}$. This quantity can be evaluated using the proximal point iteration, which is often efficient [28]. Note that running an online algorithm on smooth surrogates $\{\ell^{1/\gamma}_{t}\}_{t=1}^{T}$ differs from the existing implicit online learning (proximal point) approach [19], since proximal point is applied not as an online algorithm itself, but instead as a subroutine to find the envelope gradient $\nabla\ell_{t}^{1/\gamma}$. This smoothing strategy can be applied jointly with any existing adaptive online algorithm, such as AdaGrad.
5 Illustrative experiments
In this section, we provide toy numerical experiments to demonstrate the tightness of the $G^{\star}$ regret bound.
5.1 Experiment setup
We consider two kinds of loss functions from Section˜2.
  • regression with $\ell_{p}$ loss $\ell_{t}^{p}(x)\coloneqq\tfrac{1}{p}|\langle a_{t},x\rangle-b_{t}|^{p}$, and
  • binary classification with cross-entropy loss $\ell_{t}^{\mathrm{ce}}(x)\coloneqq\log(1+\exp(y_{t}\langle a_{t},x\rangle))$
Data preparation.
We fix $n=2$ for all the experiments. For $\ell_{p}$-regression, we let $p=4$ and generate $a_{t}\sim\mathcal{N}(0,\frac{1}{10}I_{n})$ from standard normal distribution and generate $\bar{x}\sim\mathcal{N}(0,\frac{1}{10}I_{n})$. Then we let $b_{t}=\langle a_{t},\bar{x}\rangle+\sigma\varepsilon_{t}$ with $\varepsilon_{t}\sim\mathcal{N}(0,1)$ and $\sigma\in[0.1,1]$ parametrizes the noise level. As $\sigma\rightarrow 0$, $\ell_{t}^{p}(\bar{x})\rightarrow 0$ for all $t\in[T]$. For binary classification, we generate i.i.d. $a_{t}\sim\mathcal{U}[0,1]^{n}$ from uniform distribution and $y_{t}\sim\operatorname{Bernoulli}(\delta),\delta\in[0.9,1]$. As $\delta\rightarrow 1$, it is more likely to find $\bar{x}$ such that $y_{t}\langle a_{t},\bar{x}\rangle\rightarrow$-$\infty$ and $\ell_{t}^{\mathrm{ce}}(x)\rightarrow 0$ for most of $t\in[T]$.
Algorithm configuration.
After generating $\{\ell_{t}\}_{t=1}^{T}$, we compute $x^{\star}\in\arg\min_{x\in\{x:\|x\|\leq 10^{-2}\}}$ and obtain $L_{T}^{\star},G_{T}^{\star}$ using CVX [15]. Finally, we run AdaGrad-Norm with parameter $\alpha=\frac{\sqrt{2}D}{2}$ and obtain $\rho_{T}(x^{\star})$. The range of time horizon is taken to be $T\in[10^{2},10^{4}]$.
5.2 Regret comparison
Figure˜1 plots 1) the true regret of AdaGrad-Norm; 2) the $G^{\star}_{T}$ regret upper bound predicted by Theorem˜2.2; 3) the regret upper bound predicted by Theorem˜2.2 but with $G^{\star}_{T}$ replaced by $2LL^{\star}_{T}$. The experiment shows that the $G^{\star}$ regret often yields a substantially tighter characterization of the true algorithm performance, several magnitudes tighter than the $L^{\star}$-based bound, confirming our theoretical claims.
Figure 1: Comparison between true regret and the theoretical regret upper bounds given by $L_{T}^{\star}$ and $G_{T}^{\star}$. First row: experiment on $\ell_{p}$ regression. From left to right: $T\in\{100,500,5000,10000\}$. Second row: experiment on logistic regression. From left to right: $T\in\{100,500,5000,10000\}$.
6 Conclusion
This paper proposes the $G^{\star}$ regret, a problem-dependent regret that refines the existing $L^{\star}$ regret. We establish regret upper and lower bounds and extend the $G^{\star}$ regret to other OCO settings. We also showcase the applications of the $G^{\star}$ regret in stochastic optimization. Experiments validate the efficiency of the $G^{\star}$ regret.
Acknowledgement
We appreciate the constructive feedback from Professor Qi Deng from Shanghai Jiao Tong University.
Appendix
Appendix A Proof of results in Section 2
A.1 Discussion on Assumption A1
It is known that $L$-smoothness implies A1 if $\operatornamewithlimits{dom}\ell_{t}=\mathbb{R}^{n}$. Otherwise, A1 only holds locally for points away from the boundary of $\operatornamewithlimits{dom}\ell_{t}$ [8].
Proposition A.1 ([8], Theorem 3.1).
[…]
$\ell_{t}(x)-\ell_{t}(y)-\langle\nabla\ell_{t}(y),x-y\rangle\geq\tfrac{1}{2L}\|\nabla\ell_{t}(x)-\nabla\ell_{t}(y)\|^{2}.$
According to Proposition˜A.1, it suffices to take $\mathcal{X}\subsetneq\operatornamewithlimits{dom}\ell_{t}$ such that for all $x,y\in\mathcal{X}$,
[…]
A.2 Algorithms adapting to the $G^{\star}$ regret
This section records online algorithms adaptive to the $G^{\star}$ regret that are used in the main results.
A.2.1 Online gradient descent (OGD)
The algorithm is presented in Algorithm˜1.
input initial point $x^{1}$, learning rate $\eta>0$
for $t=1,2,\dots$ do
   Play $x^{t}$ and receive $\nabla\ell_{t}(x^{t})$
  Update $x^{t+1}=\Pi_{\mathcal{X}}[x^{t}-\eta\nabla\ell_{t}(x^{t})]$
end for
Algorithm 1 Online gradient descent (OGD)
A.2.2 AdaGrad-Norm
The algorithm is presented in Algorithm˜2.
input initial point $x^{1}$, learning rate $\alpha>0$
for $t=1,2,\dots$ do
   Play $x^{t}$ and receive $\nabla\ell_{t}(x^{t})$
  Update $x^{t+1}=\Pi_{\mathcal{X}}\Big[x^{t}-\frac{\alpha}{\sqrt{\sum_{i=1}^{t}\|\nabla\ell_{i}(x^{i})\|^{2}}}\nabla\ell_{t}(x^{t})\Big]$ and skip if $\nabla\ell_{t}(x^{t})=0$
end for
Algorithm 2 AdaGrad-Norm
A.2.3 AdaFTRL
We refer the interested readers to [26] for more details of AdaFTRL. Define $h(y)=(r(x)+\mathbb{I}\{x\in\mathcal{X}\})^{\ast}$, $V_{h}(x,y)\coloneqq h(x)-h(y)-\langle\nabla h(y),x-y\rangle$ and $L_{t}\coloneqq\sum_{j=1}^{t}\nabla\ell_{j}(x^{j})$. Then recursively define
$\Delta_{t}\coloneqq\sum_{j=1}^{t}\Delta_{j-1}V_{h}(-\tfrac{L_{j}}{\Delta_{j-1}},-\tfrac{L_{j-1}}{\Delta_{i-1}})$
and let $\Delta_{j-1}V_{h}(-\tfrac{L_{j}}{\Delta_{j-1}},-\tfrac{L_{j-1}}{\Delta_{j-1}})=\lim_{a\rightarrow 0^{+}}aV_{h}(-\tfrac{L_{j}}{a},-\tfrac{L_{j-1}}{a})$ if $\Delta_{j-1}=0$. Finally, define the regularizer
$R_{t}(x)\coloneqq\Delta_{t}\cdot r(x).$
The algorithm is given in Algorithm˜3.
input initial point $x^{0}$, $\lambda$-strongly convex regularizer $r$
for $t=1,2,\dots$ do
   Play $x^{t}=\operatornamewithlimits{arg\,min}_{x\in\mathcal{X}}\{\langle L_{t-1},x\rangle+R_{t}(x)\}$ and receive $\nabla\ell_{t}(x^{t})$
end for
Algorithm 3 AdaFTRL
A.3 Auxiliary results
Lemma˜A.1 is a well-known result used in the proof of adaptive online algorithms.
Lemma A.1 ([25], Lemma 4.13).
Let $\{a_{t}\}_{t=0}^{T}$ be a sequence of non-negative numbers and $g$ be a non-increasing continuous function, then $\sum_{t=1}^{T}a_{t}g(a_{0}+\sum_{j=1}^{t}a_{j})\leq\int_{a_{0}}^{\sum_{t=1}^{T}a_{t}}g(x)~\mathrm{d}x$.
The following lemmas present the regret guarantees for AdaGrad-Norm and AdaFTRL applied to linear losses.
Theorem A.1 ([25], Theorem 4.13).
Suppose A2 holds. Then the regret of AdaGrad-Norm (Algorithm˜2) with $\alpha>0$ applied to linear losses $\{g_{t}\}_{t=1}^{T}$ satisfies $\sum_{t=1}^{T}\langle g_{t},x^{t}-x\rangle\leq(\alpha+\frac{D^{2}}{2\alpha})\sqrt{\sum_{t=1}^{T}\|g_{t}\|^{2}}$ for any $x\in\mathcal{X}$.
Theorem A.2 ([26], Theorem 1).
Suppose A2 holds and that $r:\mathcal{X}\rightarrow\mathbb{R}_{+}$ is a non-negative lower semicontinuous function that is $\lambda$-strongly convex with respect to Euclidean norm $\|\cdot\|$ and is bounded from above. Then the regret of AdaFTRL (Algorithm˜3) applied to linear losses $\{g_{t}\}_{t=1}^{T}$ satisfies $\sum_{t=1}^{T}\langle g_{t},x^{t}-x\rangle\leq\sqrt{3}\max\{D,\tfrac{1}{\sqrt{2\lambda}}\}\sqrt{\sum_{t=1}^{T}\|g_{t}\|^{2}}(1+r(x))$ for any $x\in\mathcal{X}$.
A.4 Proof of Proposition 2.1
We prove different cases separately.
Case 1.
$G_{T}^{\star}\leq 2LL^{\star}_{T}$. To show this inequality, let $x^{\star}\in\operatornamewithlimits{arg\,min}_{x\in\mathcal{X}}\sum_{t=1}^{T}\ell_{t}(x)$. Using (5),
$G^{\star}_{T}=\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{\star})\|^{2}\leq 2L\sum_{t=1}^{T}[\ell_{t}(x^{\star})-\inf\ell_{t}]=2LL^{\star}_{T},$
and this completes the proof.
Case 2.
$G_{T}^{\star}\ll L^{\star}_{T}$. To show this inequality, we adopt Example˜2.4 with $p=\tfrac{1}{6}$ and
$G_{T}^{\star}=\Theta(T^{1/3})\ll\Theta(T^{2/3})=L_{T}^{\star}.$
Case 3.
$V_{T}\ll L^{\star}_{T}$ and $V_{T}\ll G^{\star}_{T}$. To show this inequality, we adopt the example from Section 6.3 of [39] and consider the following 1D OCO problem with losses $\ell^{B}_{t}(x)=\tfrac{1}{2}(a_{t}x-1)^{2}$ and $\mathcal{X}=[-1,1]$, where $a_{t}=\tfrac{1}{2}-\tfrac{t-1}{T}$ and $T=2K+1$ for some integer $K\geq 3$. According to Instance 1 of [39],we have $V_{T}\leq\sum_{t=1}^{T}\tfrac{4}{T^{2}}\leq\tfrac{4}{T}=\mathcal{O}(1)$. Next, we consider $G_{T}^{\star}$ and $L_{T}^{\star}$. By definition, we have $\inf\ell^{B}_{t}=0$ and that
$\sum_{t=1}^{T}\ell^{B}_{t}(x)=\frac{1}{2}\sum_{t=1}^{T}(a_{t}^{2}x^{2}-2a_{t}x+1)=\tfrac{1}{2}(\sum_{t=1}^{T}a_{t}^{2})x^{2}-(\sum_{t=1}^{T}a_{t})x+\tfrac{T}{2}$
and the minimizer $x^{\star}=\tfrac{\sum_{t=1}^{T}a_{t}}{\sum_{t=1}^{T}a_{t}^{2}}=\tfrac{\sum_{t=1}^{T}\frac{1}{2}-\frac{t-1}{T}}{\sum_{t=1}^{T}(\frac{1}{2}-\frac{t-1}{T})^{2}}=\frac{6T}{T^{2}+2}\in\mathcal{X}$ gives
$L_{T}^{\star}=$ $\textstyle\sum_{t=1}^{T}\ell^{B}_{t}(x^{\star})=\tfrac{T(T^{2}-1)}{2(T^{2}+2)}=\Theta(T)$
$G_{T}^{\star}=$ $\textstyle\sum_{t=1}^{T}\|\nabla\ell^{B}_{t}(x^{\star})\|^{2}=\tfrac{-32+60T^{2}-33T^{4}+5T^{6}}{60T(2+T^{2})^{2}}=\Theta(T),$
where $G_{T}^{\star}=\Theta(T)$ can be verified by noticing that $\lim_{T\rightarrow\infty}\frac{G_{T}^{\star}}{T}=12$.
Case 4.
$L_{T}^{\star}\ll V_{T}$ and $G_{T}^{\star}\ll V_{T}$. To show this inequality, we again adopt the example from Section 6.3 of [39] and consider the following 1D OCO problem with losses $\ell^{C}_{t}(x)=\tfrac{1}{2}(a_{t}x-b_{t})^{2}$ and $\mathcal{X}=[-1,1]$, where $(a_{t},b_{t})=(1,1)$ for even $t$ and $(a_{t},b_{t})=(\tfrac{1}{2},\tfrac{1}{2})$ otherwise. According to Instance 2 of [39],we have $V_{T}=\Theta(T)$. On the other hand, $x^{\star}=1$ minimizes each $\ell^{C}_{t}$ and $G_{T}^{\star}=L_{T}^{\star}=0$. This completes the proof.
A.5 Proof of Lemma 2.1
Define $x\coloneqq\sqrt{\sum_{t=1}^{T}\|a_{t}\|^{2}}$ and $y\coloneqq\sqrt{\sum_{t=1}^{T}\|b_{t}\|^{2}}$. By Cauchy’s inequality, we have
$\sum_{t=1}^{T}\|a_{t}-b_{t}\|^{2}\geq\sum_{t=1}^{T}\|a_{t}\|^{2}+\sum_{t=1}^{T}\|b_{t}\|^{2}-2\sqrt{\sum_{t=1}^{T}\|a_{t}\|^{2}}\sqrt{\sum_{t=1}^{T}\|b_{t}\|^{2}}=(x-y)^{2}.$
Then we deduce that
$\textstyle\alpha\sqrt{\sum_{t=1}^{T}\|a_{t}\|^{2}}-\beta\sum_{t=1}^{T}\|a_{t}-b_{t}\|^{2}\leq{}$ $\alpha x-\beta(x-y)^{2}$
$\leq{}$ $\max_{x}~\{\alpha x-\beta(x-y)^{2}\}$
$={}$ $\alpha y+\tfrac{\alpha^{2}}{4\beta}$
$={}$ $\textstyle\alpha\sqrt{\sum_{t=1}^{T}\|b_{t}\|^{2}}+\tfrac{\alpha^{2}}{4\beta},$
and this completes the proof.
A.6 Proof of Theorem 2.2
Using Theorem˜A.1 with $\alpha=\frac{\sqrt{2}D}{2}$and A2, we deduce that
$\rho_{T}(x)={}$ $\textstyle\sum_{t=1}^{T}\ell_{t}(x^{t})-\ell_{t}(x)$
$\leq{}$ $\textstyle\sum_{t=1}^{T}\langle\nabla\ell_{t}(x^{t}),x^{t}-x\rangle-\tfrac{1}{2L}\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t})-\nabla\ell_{t}(x)\|^{2}$ (18)
$\leq{}$ $\textstyle\sqrt{2}D\sqrt{\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t})\|^{2}}-\tfrac{1}{2L}\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t})-\nabla\ell_{t}(x)\|^{2}$ (19)
$\leq{}$ $\textstyle\sqrt{2}D\sqrt{\sum_{t=1}^{T}\|\nabla\ell_{t}(x)\|^{2}}+LD^{2},$ (20)
where (18) applies A1, (19) uses Theorem˜A.1, and (20) uses Lemma˜2.1 with $a_{t}=\nabla\ell_{t}(x^{t}),b_{t}=\nabla\ell_{t}(x),\alpha=\sqrt{2}D$, and $\beta=\frac{1}{2L}$. This completes the proof.
A.7 Proof of Corollary 2.1
We start from the inequality (11)
$\rho_{T}(x)\leq\tfrac{\|x^{1}-x\|^{2}}{2\eta}+\sum_{t=1}^{T}[-\tfrac{1}{2L}\|\nabla\ell_{t}(x^{t})-\nabla\ell_{t}(x)\|^{2}+\tfrac{\eta}{2}\|\nabla\ell_{t}(x^{t})\|^{2}].$
Then we lower bound $\rho_{T}(x)$ as follows:
$\rho_{T}(x)={}$ $\textstyle\sum_{t=1}^{T}[\ell_{t}(x^{t})-\ell_{t}(x)]$
$={}$ $\textstyle\sum_{t=1}^{T}[\ell_{t}(x^{t})-\inf\ell_{t}+\inf\ell_{t}-\ell_{t}(x)]$
$\geq{}$ $\textstyle\tfrac{1}{2L}\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t})\|^{2}-\sum_{t=1}^{T}[\ell_{t}(x)-\inf\ell_{t}]$
$={}$ $\textstyle\tfrac{1}{2L}\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t})\|^{2}-L_{T}(x).$
Since $\eta\in[0,\tfrac{2}{L})$, we have $\tfrac{\eta L}{2}\in[0,1)$. Subtracting $\tfrac{\eta L}{2}\rho_{T}(x)$ off both sides, we deduce that
$(1-\tfrac{\eta L}{2})\rho_{T}(x)$
$\leq{}$ $\textstyle\tfrac{\|x^{1}-x\|^{2}}{2\eta}+\sum_{t=1}^{T}[-\tfrac{1}{2L}\|\nabla\ell_{t}(x^{t})-\nabla\ell_{t}(x)\|^{2}+\tfrac{\eta}{2}\|\nabla\ell_{t}(x^{t})\|^{2}]-\tfrac{\eta L}{2}\rho_{T}(x)$
$\leq{}$ $\textstyle\tfrac{\|x^{1}-x\|^{2}}{2\eta}+\sum_{t=1}^{T}[-\tfrac{1}{2L}\|\nabla\ell_{t}(x^{t})-\nabla\ell_{t}(x)\|^{2}+\tfrac{\eta}{2}\|\nabla\ell_{t}(x^{t})\|^{2}]-\tfrac{\eta}{4}\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t})\|^{2}+\tfrac{\eta L}{2}L_{T}(x)$
$={}$ $\textstyle\tfrac{\|x^{1}-x\|^{2}}{2\eta}+\tfrac{\eta L}{2}L_{T}(x)+\tfrac{1}{2}\sum_{t=1}^{T}[\underbrace{-\tfrac{1}{L}\|\nabla\ell_{t}(x^{t})-\nabla\ell_{t}(x)\|^{2}+\tfrac{\eta}{2}\|\nabla\ell_{t}(x^{t})\|^{2}}_{\heartsuit}]$
Next, for $\eta\in(0,\tfrac{2}{L})$, we further bound $\heartsuit$ by completing the squares:
$-\tfrac{1}{L}\|\nabla\ell_{t}(x^{t})-\nabla\ell_{t}(x)\|^{2}+\tfrac{\eta}{2}\|\nabla\ell_{t}(x^{t})\|^{2}$
$={}$ $(\tfrac{\eta}{2}-\tfrac{1}{L})\|\nabla\ell_{t}(x^{t})-\tfrac{1}{2-\eta L}\nabla\ell_{t}(x)\|^{2}+\tfrac{\eta}{2(2-\eta L)}\|\nabla\ell_{t}(x)\|^{2}\leq\tfrac{\eta}{2(2-\eta L)}\|\nabla\ell_{t}(x)\|^{2}$
Plugging the relation back, we deduce that
$(1-\tfrac{\eta L}{2})\rho_{T}(x)\leq\tfrac{\|x^{1}-x\|^{2}}{2\eta}+\tfrac{\eta}{2}LL_{T}(x)+\tfrac{\eta}{2(2-\eta L)}G_{T}(x).$
Dividing both sides by $1-\tfrac{\eta L}{2}$, we have $\rho_{T}(x)\leq\tfrac{\|x^{1}-x\|^{2}}{(2-\eta L)\eta}+\tfrac{\eta}{2-\eta L}[LL_{T}(x)+\tfrac{1}{2-\eta L}G_{T}(x)]$ and completes the proof.
A.8 Proof of Theorem 2.3
Recall that $\hat{\ell}_{t}(x)\coloneqq\langle\nabla\ell_{t}(x^{t}),x\rangle$. Consider the online linear optimization problem minimizing $\sum_{t=1}^{T}\hat{\ell}_{t}(x^{t})-\hat{\ell}_{t}(x)$ for any fixed decision $x\in\mathcal{X}$ in hindsight. The choice $\lambda\geq\frac{1}{2D^{2}}$ ensures that $\max\{D,\tfrac{1}{\sqrt{2\lambda}}\}=D$, and with Theorem˜A.2, AdaFTRL satisfies
$\sum_{t=1}^{T}\hat{\ell}_{t}(x^{t})-\hat{\ell}_{t}(x)\leq\sqrt{3}D\sqrt{\sum_{t=1}^{T}\|\nabla\hat{\ell}_{t}(x^{t})\|^{2}}(1+r(x))\leq\sqrt{3}RD\sqrt{\sum_{t=1}^{T}\|\nabla\hat{\ell}_{t}(x^{t})\|^{2}},$ (21)
where the last inequality applies the definition $R\coloneqq\max_{x\in\mathcal{X}}r(x)+1$. Using A2 and (9), we deduce that
$\rho_{T}(x)={}$ $\textstyle\sum_{t=1}^{T}\ell_{t}(x^{t})-\ell_{t}(x)$
$\leq{}$ $\textstyle\sum_{t=1}^{T}[\langle\nabla\ell_{t}(x^{t}),x^{t}-x\rangle-\tfrac{1}{2L}\|\nabla\ell_{t}(x^{t})-\nabla\ell_{t}(x)\|^{2}]$
$={}$ $\textstyle\sum_{t=1}^{T}[\hat{\ell}_{t}(x^{t})-\hat{\ell}_{t}(x)-\tfrac{1}{2L}\|\nabla\ell_{t}(x^{t})-\nabla\ell_{t}(x)\|^{2}]$
$\leq{}$ $\textstyle\sqrt{3}RD\sqrt{\sum_{t=1}^{T}\|\nabla\hat{\ell}_{t}(x^{t})\|^{2}}-\tfrac{1}{2L}\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t})-\nabla\ell_{t}(x)\|^{2}$ (22)
$={}$ $\textstyle\sqrt{3}RD\sqrt{\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t})\|^{2}}-\tfrac{1}{2L}\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t})-\nabla\ell_{t}(x)\|^{2},$ (23)
$\leq{}$ $\textstyle\sqrt{3}RD\sqrt{\sum_{t=1}^{T}\|\nabla\ell_{t}(x)\|^{2}}+2R^{2}LD^{2},$ (24)
where (22) plugs in (21); (23) uses the fact $\nabla\hat{\ell}_{t}(x^{t})=\nabla\ell_{t}(x^{t})$ and (24) again applies Lemma˜2.1 with $a_{t}=\nabla\ell_{t}(x^{t}),b_{t}=\nabla\ell_{t}(x),\alpha=\sqrt{3}RD$, and $\beta=\frac{1}{2L}$. This completes the proof.
A.9 Proof of Theorem 2.4
The proof follows from the standard online linear optimization lower bound construction using Rademacher random variables. According to Theorem 5.1 of [25], there exist vectors $\{g_{t}\}_{t=1}^{T}$ with $\|g_{t}\|=M$ such that
$\rho_{T}(x)=\sum_{t=1}^{T}\langle g_{t},x^{t}\rangle-\sum_{t=1}^{T}\langle g_{t},x\rangle\geq\frac{DM\sqrt{T}}{4}$
for any $x\in\mathcal{X}$. With $\ell_{t}(x)=\langle g_{t},x\rangle$, $\nabla\ell_{t}(x)=g_{t}$ is independent of $x$ and $G_{T}(x)=\sum_{t=1}^{T}\|\nabla\ell_{t}(x)\|^{2}=M^{2}T$ regardless of $x$. Hence $G_{T}^{\star}=M^{2}T$ and $\rho_{T}(x)\geq\tfrac{DM\sqrt{T}}{4}=\tfrac{D}{4}\sqrt{G_{T}^{\star}}$, which completes the proof.
Appendix B Proof of results in Section 3
B.1 Proof of Theorem 3.1
We adopt a slightly different argument compared to Theorem˜2.1. First, we deduce that
$\|x^{t+1}-\hat{x}_{t}\|^{2}\leq{}$ $\|x^{t}-\eta\nabla\ell_{t}(x^{t})-\hat{x}_{t}\|^{2}$
$={}$ $\|x^{t}-\hat{x}_{t}\|^{2}-2\eta\langle\nabla\ell_{t}(x^{t}),x^{t}-\hat{x}_{t}\rangle+\eta^{2}\|\nabla\ell_{t}(x^{t})\|^{2}$
$\leq{}$ $\|x^{t}-\hat{x}_{t}\|^{2}-2\eta[\ell_{t}(x^{t})-\ell_{t}(\hat{x}_{t})]\underbrace{-\tfrac{\eta}{L}\|\nabla\ell_{t}(x^{t})-\nabla\ell_{t}(\hat{x}_{t})\|^{2}+\eta^{2}\|\nabla\ell_{t}(x^{t})\|^{2}}_{{\heartsuit}}.$
Next, we upper bound $\heartsuit$ for $\eta\in(0,\tfrac{1}{L})$ with Young’s inequality $\|a+b\|^{2}\leq(1+\gamma)\|a\|^{2}+(1+\gamma^{-1})\|b\|^{2}$:
$-\tfrac{\eta}{L}\|\nabla\ell_{t}(x^{t})-\nabla\ell_{t}(\hat{x}_{t})\|^{2}+\eta^{2}\|\nabla\ell_{t}(x^{t})\|^{2}$
$={}$ $-\tfrac{\eta}{L}\|\nabla\ell_{t}(x^{t})-\nabla\ell_{t}(\hat{x}_{t})\|^{2}+\eta^{2}\|\nabla\ell_{t}(x^{t})-\nabla\ell_{t}(\hat{x}_{t})+\nabla\ell_{t}(\hat{x}_{t})\|^{2}$
$\leq{}$ $\eta[(1+\gamma)\eta-\tfrac{1}{L}]\|\nabla\ell_{t}(x^{t})-\nabla\ell_{t}(\hat{x}_{t})\|^{2}+(1+\gamma^{-1})\eta^{2}\|\nabla\ell_{t}(\hat{x}_{t})\|^{2}.$ (25)
Taking $\gamma=1$ and $\eta\in(0,\frac{1}{2L}]$ gives $(1+\gamma)\eta-\tfrac{1}{L}\leq 0$ and that
$\|x^{t+1}-\hat{x}_{t}\|^{2}\leq\|x^{t}-\hat{x}_{t}\|^{2}-2\eta[\ell_{t}(x^{t})-\ell_{t}(\hat{x}_{t})]+2\eta^{2}\|\nabla\ell_{t}(\hat{x}_{t})\|^{2},$
and the standard dynamic regret analysis (e.g. [38]) applies:
$\textstyle\sum_{t=1}^{T}\ell_{t}(x^{t})-\ell_{t}(\hat{x}_{t})={}$ $\textstyle\tfrac{1}{2\eta}\sum_{t=1}^{T}[\|x^{t}-\hat{x}_{t}\|^{2}-\|x^{t+1}-\hat{x}_{t}\|^{2}]+\eta\|\nabla\ell_{t}(\hat{x}_{t})\|^{2}$
$\leq{}$ $\textstyle\tfrac{D^{2}}{2\eta}+\tfrac{1}{2\eta}\sum_{t=2}^{T}[\|x^{t}-\hat{x}_{t}\|^{2}-\|x^{t}-\hat{x}_{t-1}\|^{2}]+\eta\hat{G}_{T}$
$\leq{}$ $\textstyle\tfrac{D^{2}}{2\eta}+\tfrac{D}{\eta}\sum_{t=2}^{T}\|\hat{x}_{t}-\hat{x}_{t-1}\|+\eta\hat{G}_{T}$
$={}$ $\tfrac{D(D+2\mathcal{P}_{T})}{2\eta}+\eta\hat{G}_{T}.$
This completes the proof by doing a case analysis on the value of $\eta=\min~\{\sqrt{\tfrac{{D(D+2\mathcal{P}_{T}})}{{2\hat{G}_{T}}}},\frac{1}{2L}\}$.
B.2 Dynamic regret and proof of rest of the results in Section 3.1
This section establishes an $\mathcal{O}(\sqrt{(1+2\mathcal{P}_{T}+\hat{G}_{T})(1+2\mathcal{P}_{T})}\log\log T)$ regret upper bound using the meta-expert algorithm framework developed in [39, 38]. For brevity, this paper focuses on $\mathrm{Sword}_{\mathrm{small}}$, a simpler version of the framework developed in [38]. Our analyses are compatible with the more sophisticated version in [39]. Following the setup in [38], our analysis resorts to an extra constant
$M\coloneqq\max_{t\in[T]}~\max_{x\in\mathcal{X}}\|\nabla\ell_{t}(x)\|,$
which can be further bounded by taking any $\bar{x}\in\mathcal{X}$ and
$\max_{t\in[T]}\max_{x\in\mathcal{X}}\|\nabla\ell_{t}(x)\|\leq\max_{t\in[T]}\|\nabla\ell_{t}(\bar{x})\|+LD.$
This constant $M$ will appear in a $\log\log$ term in the final result.
B.2.1 Algorithm design
$\mathrm{Sword}_{\mathrm{small}}$ adopts VanillaHedge (exponentiated gradient) of a meta algorithm to choose the learning rate for the lower-level expert algorithm (OGD, Algorithm˜1). The meta algorithm operates on a grid of learning rates defined as follows
$\mathcal{N}=\{\eta:\eta=2^{k}\cdot\eta_{\min},k\in[N]\}$
where
$\eta_{\min}\coloneqq\tfrac{D}{M}\sqrt{\tfrac{1}{2T}}\qquad\text{and}\qquad N\coloneqq\Big\lceil\tfrac{1}{\log 2}\log(\tfrac{M\sqrt{2T}}{LD})\Big\rceil.$ (26)
Then, the exponentiated gradient combines the iterates from $N$ copies of OGD with different learning rates. The whole algorithm is presented in Algorithm˜4.
input initial vector $p^{1}=\tfrac{1}{N}\textbf{1}_{N}$ and $\eta_{i}=\eta_{\min}2^{i-1},i\in[N]$, exponentiated gradient learning rate $\delta>0$, initial point $x^{1}_{i}=x^{1},i\in[N]$.
for $t=1,2,\dots$ do
   Play $x^{t}=\sum_{i=1}^{N}p_{i}^{t}x_{i}^{t}$ and receive $\nabla\ell_{t}(x^{t})$.
  Update $x^{t+1}_{i}=\Pi_{\mathcal{X}}[x^{t}_{i}-\eta_{i}\nabla\ell_{t}(x^{t})],i\in[N]$
   Update $p^{t+1}_{i}\propto p^{t}_{i}\exp(-\delta\langle\nabla\ell_{t}(x^{t}),x^{t}_{i}\rangle),i\in[N]$
end for
Algorithm 4 $\mathrm{Sword}_{\mathrm{small}}$ [38]
B.2.2 Auxiliary results
Before we proceed, we present several auxiliary results that tighten the existing analyses through A1.
Lemma B.1 (Meta-regret).
Under A1 and A2, Algorithm˜4 satisfies
$\sum_{t=1}^{T}\ell_{t}(x^{t})-\sum_{t=1}^{T}\ell_{t}(x^{t}_{i})\leq\textstyle\tfrac{2+\log N}{\delta}+\delta D^{2}\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t})\|^{2}-\tfrac{1}{2L}\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t})-\nabla\ell_{t}(x^{t}_{i})\|^{2}$
for any $i\in[N]$, where $N$ is defined in (26).
Proof.
By A1, we have, for any $i\in[N]$, that
$\textstyle\sum_{t=1}^{T}\ell_{t}(x^{t})-\sum_{t=1}^{T}\ell_{t}(x^{t}_{i})\leq\sum_{t=1}^{T}\langle\nabla\ell_{t}(x^{t}),x^{t}-x^{t}_{i}\rangle-\tfrac{1}{2L}\|\nabla\ell_{t}(x^{t})-\nabla\ell_{t}(x^{t}_{i})\|^{2}.$ (27)
The following proof is adapted from the proof of Theorem 11 of [38]. Using the regret guarantees of exponentiated gradient applied to linearized loss $\langle\nabla\ell_{t}(x^{t}),x\rangle$, we have, by Theorem 19 of [31], that,
$\textstyle\sum_{t=1}^{T}\langle\nabla\ell_{t}(x^{t}),x^{t}-x^{t}_{i}\rangle\leq{}$ $\textstyle\tfrac{2+\log N}{\delta}+\delta\sum_{t=1}^{T}\max_{i\in[N]}\langle\nabla\ell_{t}(x^{t}),x_{i}^{t}\rangle^{2}$
$\leq{}$ $\textstyle\tfrac{2+\log N}{\delta}+\delta D^{2}\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t})\|^{2}$ (28)
combining (27) with (28) completes the proof. ∎
The next result is a strengthened version of Theorem˜3.1.
Lemma B.2 (Expert regret).
Under the same conditions as Theorem˜3.1 and suppose $\eta\in(0,\tfrac{1}{4L}]$. Then
$\hat{\rho}(\{\hat{x}_{t}\}_{t=1}^{T})\leq\tfrac{D(D+2\mathcal{P}_{T})}{2\eta}+\eta\hat{G}_{T}-\tfrac{1}{4L}\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t}_{i})-\nabla\ell_{t}(\hat{x}_{t})\|^{2}.$
for all $i\in[N]$.
Proof.
For each $i\in[N]$, by (25), we have $\eta(1+\gamma)-\tfrac{1}{L}=2\eta-\tfrac{1}{L}\leq-\tfrac{1}{2L}$ when $\eta\leq\tfrac{1}{4L}$ and $\gamma=1$. Hence
$\|x^{t+1}-\hat{x}_{t}\|^{2}\leq\|x^{t}-\hat{x}_{t}\|^{2}-2\eta[\ell_{t}(x^{t})-\ell_{t}(\hat{x}_{t})]+2\eta^{2}\|\nabla\ell_{t}(\hat{x}_{t})\|^{2}-\tfrac{\eta}{2L}\|\nabla\ell_{t}(x^{t}_{i})-\nabla\ell_{t}(\hat{x}_{t})\|^{2}.$
Instead of dropping $\|\nabla\ell_{t}(x^{t}_{i})-\nabla\ell_{t}(\hat{x}_{t})\|^{2}$, keep it throughout the analysis completes the proof. ∎
B.2.3 Algorithm analysis
Theorem B.1 (Formal version of Theorem˜3.2).
Under the same conditions as Theorem˜3.1 and letting $\delta=\sqrt{\tfrac{2+\log N}{D^{2}\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t})\|^{2}}}$, Algorithm˜4 satisfies
$\textstyle\hat{\rho}(\{\hat{x}_{t}\}_{t=1}^{T})\leq{}$ $8\sqrt{(3+\log N){D^{2}}+2\mathcal{P}_{T}D}+8(3+\log N)LD^{2}+8LD\mathcal{P}_{T}$
$={}$ $\mathcal{O}(\sqrt{(1+\mathcal{P}_{T}+\hat{G}_{T})(1+\mathcal{P}_{T})}\log\log T)$
for any comparator sequence $\{\hat{x}_{t}\}_{t=1}^{T},\hat{x}_{t}\in\mathcal{X}$, where $N=\big\lceil\tfrac{1}{\log 2}\log(\tfrac{M\sqrt{2T}}{LD})\big\rceil=\mathcal{O}(\log T)$ is given by (26).
Proof.
Define $\eta^{\star}\coloneqq\min\{\sqrt{\tfrac{D(D+2\mathcal{P}_{T})}{2\hat{G}_{T}}},\tfrac{1}{4L}\}$. We start by the standard case where $\eta^{\star}=\sqrt{\tfrac{D(D+2\mathcal{P}_{T})}{2\hat{G}_{T}}}$. Since $\hat{G}_{T}=\sum_{t=1}^{T}\|\nabla\ell_{t}(\hat{x}^{t})\|^{2}\leq M^{2}T$, we have
$\eta_{\min}=\tfrac{D}{M}\sqrt{\tfrac{1}{2T}}\leq\eta^{\star}\leq\tfrac{1}{4L}=\eta_{\max}$
By construction, there must exist some $k\in\mathcal{N}$ such that $\eta_{k}=2^{k}\cdot\eta_{\min}$, $\eta_{k}\leq\eta^{\star}\leq 2\eta_{k}$, and by Lemma˜B.2,
$\sum_{t=1}^{T}\ell_{t}(x^{t}_{k})-\sum_{t=1}^{T}\ell_{t}(\hat{x}_{t})\leq 4\sqrt{D(D+2\mathcal{P}_{T})\hat{G}_{T}}-\tfrac{1}{4L}\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t}_{k})-\nabla\ell_{t}(\hat{x}_{t})\|^{2}.$ (29)
Combining Lemma˜B.1 and (29), we deduce that
$\hat{\rho}(\{\hat{x}_{t}\}_{t=1}^{T})={}$ $\textstyle\sum_{t=1}^{T}\ell_{t}(x^{t})-\sum_{t=1}^{T}\ell_{t}(\hat{x}_{t})$ (30)
$={}$ $\textstyle\sum_{t=1}^{T}\ell_{t}(x^{t})-\sum_{t=1}^{T}\ell_{t}(x^{t}_{k})+\sum_{t=1}^{T}\ell_{t}(x^{t}_{k})-\sum_{t=1}^{T}\ell_{t}(\hat{x}_{t})$
$\leq{}$ $\textstyle\tfrac{2+\log N}{\delta}+\delta D^{2}\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t})\|^{2}+4\sqrt{D(D+2\mathcal{P}_{T})\hat{G}_{T}}$
$\textstyle-\tfrac{1}{2L}\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t}_{k})-\nabla\ell_{t}(\hat{x}_{t})\|^{2}-\tfrac{1}{4L}\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t})-\nabla\ell_{t}(x^{t}_{k})\|^{2}$ (31)
$\leq{}$ $\textstyle\tfrac{2+\log N}{\delta}+\delta D^{2}\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t})\|^{2}+\tfrac{D(D+2\mathcal{P}_{T})}{2\eta_{k}}+\eta_{k}\hat{G}_{T}$
$\textstyle-\tfrac{1}{4L}\sum_{t=1}^{T}[\|\nabla\ell_{t}(x^{t}_{k})-\nabla\ell_{t}(\hat{x}_{t})\|^{2}+\|\nabla\ell_{t}(x^{t})-\nabla\ell_{t}(x^{t}_{k})\|^{2}]$
$\leq{}$ $\textstyle\tfrac{2+\log N}{\delta}+\delta D^{2}\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t})\|^{2}+4\sqrt{D(D+2\mathcal{P}_{T})\hat{G}_{T}}-\tfrac{1}{8L}\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t})-\nabla\ell_{t}(\hat{x}_{t})\|^{2},$ (32)
where (31) plugs in the meta regret Lemma˜B.1 and expert regret Lemma˜B.2; (32) uses $\|a\|^{2}+\|b\|^{2}\geq\frac{1}{2}\|a+b\|^{2}$. Finally, taking $\delta=\sqrt{\tfrac{2+\log N}{D^{2}\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t})\|^{2}}}$, gives
$\hat{\rho}(\{\hat{x}_{t}\}_{t=1}^{T})\leq 2D\sqrt{(2+\log N)\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t})\|^{2}}-\tfrac{1}{8L}\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t})-\nabla\ell_{t}(\hat{x}_{t})\|^{2}+4\sqrt{D(D+2\mathcal{P}_{T})\hat{G}_{T}}.$
Invoking Lemma˜2.1, we have
$\textstyle\hat{\rho}(\{\hat{x}_{t}\}_{t=1}^{T})\leq{}$ $2D\textstyle\sqrt{(2+\log N)\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t})\|^{2}}-\tfrac{1}{8L}\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t})-\nabla\ell_{t}(\hat{x}_{t})\|^{2}+4\sqrt{D(D+2\mathcal{P}_{T})\hat{G}_{T}}$
$\leq{}$ $\textstyle(2\sqrt{(2+\log N)D^{2}}+4\sqrt{D(D+2\mathcal{P}_{T})})\sqrt{\hat{G}_{T}}+8(2+\log N)LD^{2}$
$\leq{}$ $8\sqrt{(3+\log N){D^{2}}+2\mathcal{P}_{T}D}{\textstyle\sqrt{\hat{G}_{T}}}+8(2+\log N)LD^{2},$ (33)
where (33) uses $\sqrt{x}+\sqrt{y}\leq 2\sqrt{x+y}$ and this completes the proof when $\eta^{\star}=\sqrt{\tfrac{D(D+2\mathcal{P}_{T})}{2\hat{G}_{T}}}$.
Finally, we consider the case where $\eta^{\star}=\frac{1}{4L}$, in which case $\hat{G}_{T}\leq 8L^{2}D(D+2\mathcal{P}_{T})$ and for $\eta_{k}=\eta_{\max}$, we have
$\sum_{t=1}^{T}\ell_{t}(x^{t}_{k})-\sum_{t=1}^{T}\ell_{t}(\hat{x}_{t})\leq 4LD(D+2\mathcal{P}_{T})-\tfrac{1}{4L}\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t}_{k})-\nabla\ell_{t}(\hat{x}_{t})\|^{2}.$
Repeating the same regret analysis from (30) to (32) gives
$\textstyle\hat{\rho}(\{\hat{x}_{t}\}_{t=1}^{T})\leq{}$ $2D\textstyle\sqrt{(2+\log N)\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t})\|^{2}}-\tfrac{1}{8L}\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t})-\nabla\ell_{t}(\hat{x}_{t})\|^{2}+4LD(D+2\mathcal{P}_{T})$
$\leq{}$ $\textstyle 2\sqrt{(2+\log N)D^{2}}\sqrt{\hat{G}_{T}}+8(2+\log N)LD^{2}+4LD(D+2\mathcal{P}_{T})$ (34)
Combining (33) and (34) completes the proof. ∎
Remark 8.
As mentioned in [38] the choice of $\delta$ in the meta algorithm still requires knowledge of unknown quantify $\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t})\|^{2}$. This issue can be addressed using the doubling trick or an adaptive version of the hedge algorithm discussed in Appendix B of [39] or Section 7.6 of [25].
B.3 Bandit feedback and proof of results in Section 3.2
This section studies online algorithms adaptive to the $G^{\star}$ regret under bandit feedback. Specifically, we focus on BGD with a constant learning rate and BGD with an AdaGrad-Norm learning rate. The details of both algorithms are provided below.
input initial point $x^{0}$, learning rate $\eta>0$
for $t=1,2,\dots$ do
   Draw a random direction $s_{t}$ uniformly from the unit sphere $\mathbb{S}_{n}$
  Play $(y^{t}_{+},y^{t}_{-})=(x^{t}+\mu s_{t},x^{t}-\mu s_{t})$ and receive $\ell_{t}(y^{t}_{+})$, $\ell_{t}(y^{t}_{-})$
  Construct $g_{t}=\tfrac{n}{2\mu}\bigl[\ell_{t}(y^{t}_{+})-\ell_{t}(y^{t}_{-})\bigr]s_{t}$
  Update $x^{t+1}=\Pi_{\mathcal{X}}[x^{t}-\eta g_{t}]$
end for
Algorithm 5 Bandit gradient descent (BGD)
input initial point $x^{0}$, learning rate $\alpha>0$, smoothing parameter $\mu>0$
for $t=1,2,\dots$ do
   Draw a random direction $s_{t}$ uniformly from the unit sphere $\mathbb{S}_{n}$
  Play $(y^{t}_{+},y^{t}_{-})=(x^{t}+\mu s_{t},x^{t}-\mu s_{t})$ and receive $\ell_{t}(y^{t}_{+})$, $\ell_{t}(y^{t}_{-})$
  Construct $g_{t}=\tfrac{n}{2\mu}\bigl[\ell_{t}(y^{t}_{+})-\ell_{t}(y^{t}_{-})\bigr]s_{t}$
  Update $x^{t+1}=\Pi_{\mathcal{X}}\Big[x^{t}-\frac{\alpha}{\sqrt{\sum_{i=1}^{t}\|g_{i}\|^{2}}}g_{t}\Big]$ and skip if $g_{t}=0$
end for
Algorithm 6 BGD with AdaGrad-Norm learning rate
B.3.1 Auxiliary results
Lemma˜B.3 presents properties of the smoothed function $\tilde{\ell}$. Parts of these results can be found in [11].
Lemma B.3.
Under A1 and A2, the smoothed function $\tilde{\ell}_{t}$ defined in (14) is both convex and $L$-smooth. Its gradient and function value satisfy $\|\nabla\tilde{\ell}_{t}(x)-\nabla\ell_{t}(x)\|\leq L\mu$ and $\ell_{t}(x)\leq\tilde{\ell}_{t}(x)\leq\ell_{t}(x)+\tfrac{L\mu^{2}}{2}$ for all $x\in\mathcal{X}$. Moreover, the second moment of the estimator $g_{t}$ in (13) satisfies the bound
$\mathbb{E}_{s_{t}\sim\mathbb{S}_{n}}[\|g_{t}\|^{2}|x^{t}]\leq 2n\|\nabla\ell_{t}(x^{t})\|^{2}+\frac{1}{2}n^{2}\mu^{2}L^{2}.$ (35)
B.3.2 Proof of Theorem 3.3
Using the update rule of BGD, we first observe that $\|x^{t+1}-x\|^{2}\leq\|x^{t}-x\|^{2}-2\eta\langle g_{t},x^{t}-x\rangle+\eta^{2}\|g_{t}\|^{2}$. Taking the expectation with respect to $s_{t}$ conditioned on $x^{t}$ yields
$\mathbb{E}_{s_{t}}[\|x^{t+1}-x\|^{2}|x^{t}]\leq{}$ $\|x^{t}-x\|^{2}-2\eta\langle\nabla\tilde{\ell}_{t}(x^{t}),x^{t}-x\rangle+\eta^{2}\mathbb{E}_{s_{t}}[\|g_{t}\|^{2}|x^{t}]$
$\leq{}$ $\|x^{t}-x\|^{2}-2\eta\langle\nabla\tilde{\ell}_{t}(x^{t}),x^{t}-x\rangle+\eta^{2}(2n\|\nabla\ell_{t}(x^{t})\|^{2}+\tfrac{n^{2}L^{2}\mu^{2}}{2})$
$\leq{}$ $\|x^{t}-x\|^{2}-2\eta\langle\nabla\tilde{\ell}_{t}(x^{t}),x^{t}-x\rangle+\eta^{2}(4n\|\nabla\ell_{t}(x^{t})\|^{2}+4n\|\nabla\ell_{t}(x^{t})-\nabla\tilde{\ell}_{t}(x^{t})\|^{2}+\tfrac{n^{2}L^{2}\mu^{2}}{2})$
$\leq{}$ $\|x^{t}-x\|^{2}-2\eta\langle\nabla\tilde{\ell}_{t}(x^{t}),x^{t}-x\rangle+4n\eta^{2}\|\nabla\tilde{\ell}_{t}(x^{t})\|^{2}+\eta^{2}n^{2}L^{2}\mu^{2},$ (36)
where (36) follows from Lemma˜B.3 and that $4n\leq\tfrac{n^{2}}{2}$ for $n\geq 8$. By the dual strong convexity of $\tilde{\ell}_{t}(\cdot)$, we have
$\mathbb{E}_{s_{t}}[\|x^{t+1}-x\|^{2}|x^{t}]$
$\leq{}$ $\|x^{t}-x\|^{2}-2\eta[\tilde{\ell}_{t}(x^{t})-\tilde{\ell}_{t}(x)]-\tfrac{\eta}{L}\|\nabla\tilde{\ell}_{t}(x^{t})-\nabla\tilde{\ell}_{t}(x)\|^{2}+4n\eta^{2}\|\nabla\tilde{\ell}_{t}(x^{t})\|^{2}+\eta^{2}n^{2}L^{2}\mu^{2}$
[…]
$+\tfrac{\eta}{L}(4n\eta L-1)\|\nabla\tilde{\ell}_{t}(x^{t})-\tfrac{1}{1-4n\eta L}\nabla\tilde{\ell}_{t}(x)\|^{2}$
For $\eta\in(0,\tfrac{1}{4nL})$, $4n\eta L-1<0$ and
$\mathbb{E}_{s_{t}}[\|x^{t+1}-x\|^{2}|x^{t}]\leq\|x^{t}-x\|^{2}-2\eta[\tilde{\ell}_{t}(x^{t})-\tilde{\ell}_{t}(x)]+\tfrac{4n\eta^{2}}{1-4n\eta L}\|\nabla\tilde{\ell}_{t}(x)\|^{2}+\eta^{2}n^{2}L^{2}\mu^{2}$
Taking total expectation, we have
$\mathbb{E}[\|x^{t+1}-x\|^{2}]\leq\mathbb{E}[\|x^{t}-x\|^{2}]-2\eta\mathbb{E}[\tilde{\ell}_{t}(x^{t})-\tilde{\ell}_{t}(x)]+\tfrac{4n\eta^{2}}{1-4n\eta L}\|\nabla\tilde{\ell}_{t}(x)\|^{2}+\eta^{2}n^{2}L^{2}\mu^{2}.$
A re-arrangement gives the standard regret inequality, and the regret result follows by telescoping:
$\textstyle\sum_{t=1}^{T}\mathbb{E}[\tilde{\ell}_{t}(x^{t})-\tilde{\ell}_{t}(x)]\leq\tfrac{D^{2}}{2\eta}+\tfrac{2n\eta}{1-4n\eta L}\sum_{t=1}^{T}\|\nabla\tilde{\ell}_{t}(x)\|^{2}+\tfrac{\eta T}{2}n^{2}L^{2}\mu^{2}.$
By Lemma˜B.3, we have $\ell_{t}(x)\leq\tilde{\ell}_{t}(x)\leq\ell_{t}(x)+\tfrac{L\mu^{2}}{2}$ and $\|\nabla\tilde{\ell}_{t}(x)-\nabla\ell_{t}(x)\|\leq L\mu$ for any $x\in\mathcal{X}$. Therefore, the above inequality yields
$\textstyle\sum_{t=1}^{T}\mathbb{E}[\ell_{t}(x^{t})-\ell_{t}(x)]\leq{}$ $\textstyle\tfrac{D^{2}}{2\eta}+\tfrac{4n\eta}{1-4n\eta L}\sum_{t=1}^{T}\|\nabla\ell_{t}(x)\|^{2}+\tfrac{4n\eta}{1-4n\eta L}TL^{2}\mu^{2}+\tfrac{\eta T}{2}n^{2}L^{2}\mu^{2}+\tfrac{TL\mu^{2}}{2}$
$\leq{}$ $\textstyle\tfrac{D^{2}}{2\eta}+8n\eta\sum_{t=1}^{T}\|\nabla\ell_{t}(x)\|^{2}+nTL^{2}\mu^{2},$
where we substitute $\eta\in(0,\tfrac{1}{8nL}]$ and $n\geq 8$ in the last step. In particular, choosing $x=x^{\star}$, $\eta=\min\big\{\tfrac{D}{4\sqrt{nG^{\star}_{T}}},\tfrac{1}{8nL}\big\}$ and $\mu=\tfrac{D}{\sqrt{2nLT}}$ gives $\mathbb{E}[\rho_{T}(x^{\star})]\leq\max\{8nLD^{2},4D\sqrt{nG^{\star}_{T}}\}+\tfrac{LD^{2}}{2}$ and completes the proof.
B.3.3 Proof of Theorem 3.4
First, invoking Theorem˜A.1, we have
$\textstyle\sum_{t=1}^{T}\mathbb{E}[\langle\nabla\tilde{\ell}_{t}(x^{t}),x^{t}-x\rangle]=\sum_{t=1}^{T}\mathbb{E}[\langle\mathbb{E}[g_{t}|x^{t}],x^{t}-x\rangle]\leq(\frac{D^{2}}{2\alpha}+\alpha)\mathbb{E}\Big[\sqrt{\sum_{t=1}^{T}\|g_{t}\|^{2}}\Big].$
Applying the property of dual strong convexity yields
$\textstyle\sum_{t=1}^{T}\mathbb{E}[\tilde{\ell}_{t}(x^{t})-\tilde{\ell}_{t}(x)]\leq{}$ $\textstyle(\frac{D^{2}}{2\alpha}+\alpha)\mathbb{E}\big[\sqrt{\sum_{t=1}^{T}\|g_{t}\|^{2}}\big]-\frac{1}{2L}\sum_{t=1}^{T}\mathbb{E}[\|\nabla\tilde{\ell}_{t}(x^{t})-\nabla\tilde{\ell}_{t}(x)\|^{2}]$
$\leq{}$ $\textstyle(\frac{D^{2}}{2\alpha}+\alpha)\sqrt{\sum_{t=1}^{T}\mathbb{E}[\|g_{t}\|^{2}]}-\frac{1}{2L}\sum_{t=1}^{T}\mathbb{E}[\|\nabla\tilde{\ell}_{t}(x^{t})-\nabla\tilde{\ell}_{t}(x)\|^{2}]$
$\leq{}$ $\textstyle(\frac{D^{2}}{2\alpha}+\alpha)\sqrt{\sum_{t=1}^{T}2n\mathbb{E}[\|\nabla\ell_{t}(x^{t})\|^{2}]+\frac{Tn^{2}L^{2}\mu^{2}}{2}}-\frac{1}{2L}\sum_{t=1}^{T}\mathbb{E}[\|\nabla\tilde{\ell}_{t}(x^{t})-\nabla\tilde{\ell}_{t}(x)\|^{2}],$
where we use Lemma˜B.3 in the last step. Choose $\alpha=\tfrac{\sqrt{2}D}{2}$. We conclude
$\textstyle\sum_{t=1}^{T}\mathbb{E}[\tilde{\ell}_{t}(x^{t})-\tilde{\ell}_{t}(x)]\leq{}$ $\textstyle\sqrt{2}D\sqrt{\sum_{t=1}^{T}2n\mathbb{E}[\|\nabla\ell_{t}(x^{t})\|^{2}]}-\frac{1}{2L}\sum_{t=1}^{T}\mathbb{E}[\|\nabla\tilde{\ell}_{t}(x^{t})-\nabla\tilde{\ell}_{t}(x)\|^{2}]+DnL\mu\sqrt{T}$
$\leq{}$ $\textstyle 2D\sqrt{n}\sqrt{\sum_{t=1}^{T}\mathbb{E}[\|\nabla\tilde{\ell}_{t}(x^{t})\|^{2}]}-\frac{1}{2L}\sum_{t=1}^{T}\mathbb{E}[\|\nabla\tilde{\ell}_{t}(x^{t})-\nabla\tilde{\ell}_{t}(x)\|^{2}]+2DnL\mu\sqrt{T},$
where the last step follows from $\|\nabla\tilde{\ell}_{t}(x)-\nabla\ell_{t}(x)\|\leq L\mu$ and that $2\sqrt{n}\leq n$ for $n\geq 8$, To proceed, note that
$\textstyle\mathbb{E}[\|\nabla\tilde{\ell}_{t}(x^{t})-\nabla\tilde{\ell}_{t}(x)\|^{2}]\geq(1-\gamma)\mathbb{E}[\|\nabla\tilde{\ell}_{t}(x^{t})\|^{2}]+(1-\gamma^{-1})\|\nabla\tilde{\ell}_{t}(x)\|^{2},\text{~for all~}\gamma\in(0,1],$
which implies
$\textstyle\sum_{t=1}^{T}\mathbb{E}[\tilde{\ell}_{t}(x^{t})-\tilde{\ell}_{t}(x)]$
$\leq\$$\textstyle 2D\sqrt{n}\sqrt{\sum_{t=1}^{T}\mathbb{E}[\|\nabla\tilde{\ell}_{t}(x^{t})\|^{2}]}-\frac{1-\gamma}{2L}\sum_{t=1}^{T}\mathbb{E}[\|\nabla\tilde{\ell}_{t}(x^{t})\|^{2}]+\frac{\gamma^{-1}-1}{2L}\sum_{t=1}^{T}\|\nabla\tilde{\ell}_{t}(x)\|^{2}+2DnL\mu\sqrt{T}.$
Next, we do case analysis.
Case 1.
$\sum_{t=1}^{T}\mathbb{E}[\|\nabla\tilde{\ell}_{t}(x^{t})\|^{2}]\leq 64L^{2}D^{2}n$. Choosing $\gamma=1$ gives
$\textstyle\sum_{t=1}^{T}\mathbb{E}[\tilde{\ell}_{t}(x^{t})-\tilde{\ell}_{t}(x)]\leq 2D\sqrt{n}\sqrt{\sum_{t=1}^{T}\mathbb{E}[\|\nabla\tilde{\ell}_{t}(x^{t})\|^{2}]}+2DnL\mu\sqrt{T}\leq 16LD^{2}n+2DnL\mu\sqrt{T}.$
Case 2. $\sum_{t=1}^{T}\mathbb{E}[\|\nabla\tilde{\ell}_{t}(x^{t})\|^{2}]\geq 64L^{2}D^{2}n$. Then taking $\gamma=1-\tfrac{4LD\sqrt{n}}{\sqrt{\sum_{t=1}^{T}\mathbb{E}[\|\nabla\tilde{\ell}_{t}(x^{t})\|^{2}]}}\geq\tfrac{1}{2}$ gives
$\textstyle 2D\sqrt{n}\sqrt{\sum_{t=1}^{T}\mathbb{E}[\|\nabla\tilde{\ell}_{t}(x^{t})\|^{2}]}-\frac{1-\gamma}{2L}\sum_{t=1}^{T}\mathbb{E}[\|\nabla\tilde{\ell}_{t}(x^{t})\|^{2}]=0.$
Consequently, we deduce that
$\textstyle\sum_{t=1}^{T}\mathbb{E}[\tilde{\ell}_{t}(x^{t})-\tilde{\ell}_{t}(x)]\leq\$$\textstyle\frac{\gamma^{-1}-1}{2L}\sum_{t=1}^{T}\|\nabla\tilde{\ell}_{t}(x)\|^{2}+2DnL\mu\sqrt{T}$
$=\$$\textstyle\frac{2D\sqrt{n}}{\sqrt{\sum_{t=1}^{T}\mathbb{E}[\|\nabla\tilde{\ell}_{t}(x^{t})\|^{2}]}-4LD\sqrt{n}}\sum_{t=1}^{T}\|\nabla\tilde{\ell}_{t}(x)\|^{2}+2DnL\mu\sqrt{T}$
$\leq\$$\textstyle\frac{4D\sqrt{n}}{\sqrt{\sum_{t=1}^{T}\mathbb{E}[\|\nabla\tilde{\ell}_{t}(x^{t})\|^{2}]}}\sum_{t=1}^{T}\|\nabla\tilde{\ell}_{t}(x)\|^{2}+2DnL\mu\sqrt{T},$
where the last inequality uses
$\textstyle\sqrt{\sum_{t=1}^{T}\mathbb{E}[\|\nabla\tilde{\ell}_{t}(x^{t})\|^{2}]}-4LD\sqrt{n}\geq\tfrac{1}{2}\sqrt{\sum_{t=1}^{T}\mathbb{E}[\|\nabla\tilde{\ell}_{t}(x^{t})\|^{2}]}$
since $\sqrt{\sum_{t=1}^{T}\mathbb{E}[\|\nabla\tilde{\ell}_{t}(x^{t})\|^{2}]}\geq 8LD\sqrt{n}$. Together with $\sum_{t=1}^{T}\mathbb{E}[\tilde{\ell}_{t}(x^{t})-\tilde{\ell}_{t}(x)]\leq 2D\sqrt{n}\sqrt{\sum_{t=1}^{T}\mathbb{E}[\|\nabla\tilde{\ell}_{t}(x^{t})\|^{2}]}+2DnL\mu\sqrt{T}$, applying $\min\{x,\tfrac{a}{x}\}\leq\sqrt{a}$ gives
$\textstyle\sum_{t=1}^{T}\mathbb{E}[\tilde{\ell}_{t}(x^{t})-\tilde{\ell}_{t}(x)]\leq{}$ $\textstyle 2D\sqrt{n}\min\{\sqrt{\sum_{t=1}^{T}\mathbb{E}[\|\nabla\tilde{\ell}_{t}(x^{t})\|^{2}]},\tfrac{2\sum_{t=1}^{T}\|\nabla\tilde{\ell}_{t}(x)\|^{2}}{\sqrt{\sum_{t=1}^{T}\mathbb{E}[\|\nabla\tilde{\ell}_{t}(x^{t})\|^{2}]}}\}+2DnL\mu\sqrt{T}$
$\leq\$$\textstyle 2\sqrt{2}D\sqrt{n}\sqrt{\sum_{t=1}^{T}\|\nabla\tilde{\ell}_{t}(x)\|^{2}}+2DnL\mu\sqrt{T}.$
Now, putting the two cases together gives
$\textstyle\sum_{t=1}^{T}\mathbb{E}[\tilde{\ell}_{t}(x^{t})-\tilde{\ell}_{t}(x)]\leq\max\{16LD^{2}n,2\sqrt{2}D\sqrt{n}\sqrt{\sum_{t=1}^{T}\|\nabla\tilde{\ell}_{t}(x)\|^{2}}\}+2DnL\mu\sqrt{T}.$
Finally, using the inequalities $\ell_{t}(x)\leq\tilde{\ell}_{t}(x)\leq\ell_{t}(x)+\tfrac{L\mu^{2}}{2}$ and $\|\nabla\tilde{\ell}_{t}(x)-\nabla\ell_{t}(x)\|\leq L\mu$ for any $x\in\mathcal{X}$, we have
$\textstyle\mathbb{E}[\rho_{T}(x)]\leq\max\{16LD^{2}n,4D\sqrt{nG_{T}(x)}\}+2DnL\mu\sqrt{T}(n+2\sqrt{n})+\tfrac{TL\mu^{2}}{2}.$
Recall that $2\sqrt{n}\leq n$ since $n\geq 8$. Substituting $\mu=\tfrac{D}{n\sqrt{T}}$ gives $\mathbb{E}[\rho_{T}(x)]\leq\max\{16LD^{2}n,4D\sqrt{nG_{T}(x)}\}+5LD^{2}$. The proof is completed.
Appendix C Proof of results in Section 4
C.1 Proof of Corollary 4.1
We first prove the result for OGD. According to Theorem˜2.1, for any $x\in\mathcal{X}$ and $\eta\in(0,\tfrac{1}{L})$, it holds that
$\sum_{t=1}^{T}\ell_{t}(x^{t})-\ell_{t}(x)\leq\tfrac{1}{2\eta}\|x^{1}-x\|^{2}+\tfrac{\eta}{2(1-\eta L)}\sum_{t=1}^{T}\|\nabla\ell_{t}(x)\|^{2}.$
Let $x=x^{\star}$ and $\ell_{t}(x)=f(x,\xi^{t})$. We have
$\textstyle\sum_{t=1}^{T}[f(x^{t},\xi^{t})-f(x^{\star},\xi^{t})]\leq\tfrac{1}{2\eta}\|x^{1}-x\|^{2}+\tfrac{\eta}{2(1-\eta L)}\sum_{t=1}^{T}\|\nabla f(x^{\star},\xi^{t})\|^{2}.$
Note that $\mathbb{E}_{\xi^{t}}[f(x^{t},\xi^{t})|x^{t}]=f(x^{t})$ and $\mathbb{E}_{\xi^{t}}[f(x^{\star},\xi^{t})|x^{t}]=f(x^{\star})$. Taking expectation and using convexity,
$T[f(\bar{x}^{T})-f(x^{\star})]\leq{}$ $\textstyle\mathbb{E}\big[\sum_{t=1}^{T}[f(x^{t})-f(x^{\star})]\big]$
$\leq{}$ $\textstyle\tfrac{1}{2\eta}\|x^{1}-x\|^{2}+\tfrac{\eta}{2(1-\eta L)}\sum_{t=1}^{T}\mathbb{E}[\|\nabla f(x^{\star},\xi^{t})\|^{2}]$
$={}$ $\tfrac{1}{2\eta}\|x^{1}-x\|^{2}+\tfrac{\eta}{2(1-\eta L)}T(\sigma_{g}^{\star})^{2}.$
By a similar online-to-batch argument, the corresponding results for AdaGrad-Norm and AdaFTRL can be proved by noticing that $\mathbb{E}\Big[\sqrt{\sum_{t=1}^{T}\|\nabla f(x^{\star},\xi^{t})\|^{2}}\Big]\leq\sqrt{\mathbb{E}\big[\sum_{t=1}^{T}\|\nabla f(x^{\star},\xi^{t})\|^{2}\big]}=\sigma_{g}^{\star}$.
C.2 Proof of Corollary 4.2
Combining Theorem 1 of [5] with the regret guarantees in Section˜2 and the online-to-batch argument from Corollary˜4.1 completes the proof.
C.3 Proof of Theorem 4.1
We show the result by establishing a regret version of the result with $G_{T}^{\star}=0$. i.e. $\nabla\ell_{t}(x^{\star})=0$ for all $t$. Recall the update of AdaGrad-Norm for optimization problem $x^{t+1}=x^{t}-\eta_{t}\nabla\ell_{t}(x^{t})$, where $\eta_{t}=\tfrac{\alpha}{\sqrt{\sum_{j=1}^{t}\|\nabla\ell_{t}(x^{t})\|^{2}}}$.
We deduce that
$\|x^{t+1}-x\|^{2}={}$ $\|x^{t}-\eta_{t}\nabla\ell_{t}(x^{t})-x\|^{2}$
$={}$ $\|x^{t}-x\|^{2}-2\eta_{t}\langle\nabla\ell_{t}(x^{t}),x^{t}-x\rangle+\eta_{t}^{2}\|\nabla\ell_{t}(x^{t})\|^{2}$
$\leq{}$ $\|x^{t}-x\|^{2}-2\eta_{t}[\ell_{t}(x^{t})-\ell_{t}(x)]+\tfrac{\eta_{t}^{2}}{1-\eta_{t}L}\|\nabla\ell_{t}(x)\|^{2}+(\eta^{2}_{t}-\tfrac{\eta_{t}}{L})\|\nabla\ell_{t}(x^{t})-\tfrac{1}{1-\eta_{t}L}\nabla\ell_{t}(x)\|^{2}$
Taking $x=x^{\star}$, we have $\ell_{t}(x^{t})-\ell_{t}(x^{\star})\geq 0$ and that $\nabla\ell_{t}(x^{\star})=0$, giving
$\|x^{t+1}-x^{\star}\|^{2}\leq{}$ $\textstyle\|x^{1}-x^{\star}\|^{2}+\sum_{j=1}^{t}(\eta^{2}_{j}-\tfrac{\eta_{j}}{L})\|\nabla\ell_{j}(x^{j})\|^{2}$
$\leq{}$ $\textstyle\|x^{1}-x^{\star}\|^{2}+(\eta^{2}_{1}-\frac{\eta_{1}}{L})\|\nabla\ell_{1}(x^{1})\|^{2}+\sum_{j\in\{k:k\geq 2,\eta_{k}\geq\frac{1}{L}\}}(\eta^{2}_{j}-\tfrac{\eta_{j}}{L})\|\nabla\ell_{j}(x^{j})\|^{2}$
$={}$ $\textstyle\|x^{1}-x^{\star}\|^{2}+(\alpha^{2}-\tfrac{\alpha}{L}\|\nabla\ell_{1}(x^{1})\|)+\sum_{j\in\{k:k\geq 2,\eta_{k}\geq\frac{1}{L}\}}(\eta^{2}_{j}-\tfrac{\eta_{j}}{L})\|\nabla\ell_{j}(x^{j})\|^{2}$
since $\eta^{2}-\tfrac{\eta}{L}\leq 0$ for $\eta\leq\tfrac{1}{L}$. Next we bound $\sum_{j\in\{k:k\geq 2,\eta_{k}\geq\frac{1}{L}\}}(\eta^{2}_{j}-\tfrac{\eta_{j}}{L})\|\nabla\ell_{j}(x^{j})\|^{2}$ by taking $a_{0}=\|\nabla\ell_{1}(x^{1})\|^{2}$, $a_{j}=\|\nabla\ell_{j+1}(x^{j+1})\|^{2}$ and $g(u)=\tfrac{\alpha^{2}}{u}-\frac{\alpha}{L\sqrt{u}}$ in Lemma˜A.1 (note that $g$ is non-increasing for $u\leq(2L\alpha)^{2}$, which corresponds to $\eta\geq\frac{1}{2L}$ and is satisfied for $\eta_{k}\geq\frac{1}{L}$):
$\textstyle\sum_{j\in\{k:k\geq 2,\eta_{k}\geq{}\frac{1}{L}\}}(\eta^{2}_{j}-\tfrac{\eta_{j}}{L})\|\nabla\ell_{j}(x^{j})\|^{2}\leq{}$ $\textstyle\int_{\|\nabla\ell_{1}(x^{1})\|^{2}}^{(\alpha L)^{2}}g(u)~\mathrm{d}u$
$={}$ $2\alpha^{2}\log\tfrac{\alpha L}{\|\nabla\ell_{1}(x^{1})\|}-\tfrac{2\alpha}{L}[\alpha L-\|\nabla\ell_{1}(x^{1})\|]$
$={}$ $2\alpha^{2}\log\tfrac{\alpha L}{\|\nabla\ell_{1}(x^{1})\|}-2\alpha^{2}+\tfrac{2\alpha}{L}\|\nabla\ell_{1}(x^{1})\|$
Putting things together, we have
$\|x^{t+1}-x^{\star}\|^{2}\leq{}$ $\|x^{1}-x^{\star}\|^{2}+\alpha^{2}-\tfrac{\alpha}{L}\|\nabla\ell_{1}(x^{1})\|+2\alpha^{2}\log\tfrac{\alpha L}{\|\nabla\ell_{1}(x^{1})\|}-2\alpha^{2}+\tfrac{2\alpha}{L}\|\nabla\ell_{1}(x^{1})\|$
$={}$ $\|x^{1}-x^{\star}\|^{2}+\tfrac{\alpha}{L}\|\nabla\ell_{1}(x^{1})\|+2\alpha^{2}\log\tfrac{\alpha L}{\sqrt{\mathrm{e}}\|\nabla\ell_{1}(x^{1})\|}=:{}\hat{D}$
for all $t\geq 1$. Hence, the standard convergence analysis of AdaGrad-Norm with $D=\hat{D}$ applies. By Theorem˜A.1,
$\textstyle\sum_{t=1}^{T}\ell_{t}(x^{t})-\ell_{t}(x^{\star})\leq{}$ $(\textstyle\tfrac{\hat{D}}{2\alpha}+\alpha)\sqrt{\sum_{j=1}^{t}\|\nabla\ell_{t}(x^{t})\|^{2}}-\tfrac{1}{2L}\sum_{j=1}^{t}\|\nabla\ell_{t}(x^{t})\|^{2}\leq\tfrac{L}{2}(\tfrac{\hat{D}}{2\alpha}+\alpha)^{2}.$
Plugging in the definition of $\hat{D}$ gives a convergence rate of
$\textstyle\sum_{t=1}^{T}\ell_{t}(x^{t})-\ell_{t}(x^{\star})\leq{}$ $\textstyle\tfrac{L}{2}\big(\tfrac{\|x^{1}-x^{\star}\|^{2}}{2\alpha}+\alpha\log(\tfrac{\sqrt{\mathrm{e}}\alpha L}{\|\nabla\ell_{t}(x^{1})\|})+\tfrac{1}{2L}\|\nabla\ell_{t}(x^{1})\|\big)^{2}$
and an online-to-batch argument completes the proof.
C.4 Proof of Proposition 4.1
The construction modifies the standard OCO lower bound. Consider a 1D OCO problem with
$\ell_{t}(x)=\tfrac{1+\sigma_{t}}{2}g(x)+\tfrac{1-\sigma_{t}}{2}g(-x)$
where $g(x)=\bigg\{\begin{array}[]{ll}x+\tfrac{3}{2},&x\geq-1\\ \tfrac{1}{2}[x+2]_{+}^{2},&\text{else}\end{array}$ and $\sigma_{t}\in\{-1,1\}$ is a Rademacher random variable. It is easy to check that $g$ is $L=1$-smooth, convex, with $\inf g=0$. Moreover, for $\mathcal{X}=[-\delta,\delta]\subseteq[-1,1]$, we have $\ell_{t}(x)=\sigma_{t}x+\tfrac{3}{2}$ and therefore, the regret in terms of $\{\ell_{t}\}_{t=1}^{T}$ coincides with that of the 1D OCO lower bound since for $x^{t},x\in\mathcal{X}$, we have
$\rho_{T}(x)={}$ $\textstyle\sum_{t=1}^{T}\ell_{t}(x^{t})-\sum_{t=1}^{T}\ell_{t}(x)$
$={}$ $\textstyle\sum_{t=1}^{T}[\sigma_{t}x^{t}+\tfrac{3}{2}]-\sum_{t=1}^{T}[\sigma_{t}x+\tfrac{3}{2}]$
$={}$ $\textstyle\sum_{t=1}^{T}\sigma_{t}(x^{t}-x).$
Furthermore, we can upperbound $L^{\mathcal{X}}_{T}(x)$ by
$L^{\mathcal{X}}_{T}(x)=\sum_{t=1}^{T}\ell_{t}(x)-\min_{u\in\mathcal{X}}\ell_{t}(u)=\sum_{t=1}^{T}\sigma_{t}x-\min_{u\in[-\delta,\delta]}\sigma_{t}u\leq 2\delta T.$
Suppose an $\mathcal{O}(LD^{2}+D\sqrt{LL^{\mathcal{X}}_{T}(x^{\star})})$ upper bound holds, then with $D=2\delta$, it would imply
$\rho_{T}(x)\leq\mathcal{O}(4\delta^{2}+2\delta\sqrt{2\delta T})=\mathcal{O}(\delta^{2}+\delta^{3/2}\sqrt{T}).$
However, the lower bound of OCO implies $\rho_{T}(x)\geq\Omega(D\sqrt{T})=\Omega(\delta\sqrt{T})$, and taking $\delta=T^{-\beta},\beta\in(0,\tfrac{1}{3}]$ leads to an upper bound of $\mathcal{O}(T^{\frac{1-3\beta}{2}})$ and a lower bound of $\Omega(T^{\frac{1}{2}-\beta})$ that contradict each other. Dividing the upper and lower bounds completes the proof.
C.5 Proof of Proposition 4.2
We have $G^{\mathcal{X}}_{T}(x)\leq 2\gamma[\sum_{t=1}^{T}\ell^{1/\gamma}_{t}(x)-\inf\ell^{1/\gamma}_{t}]$ due to smoothness of $\ell^{1/\gamma}(x)$. The right-hand side can be further bounded by
$\textstyle\sum_{t=1}^{T}\ell^{1/\gamma}_{t}(x)-\inf\ell^{1/\gamma}_{t}={}$ $\textstyle\sum_{t=1}^{T}\ell^{1/\gamma}_{t}(x)-\min_{u\in\mathcal{X}}\ell_{t}(u)$
$\leq{}$ $\textstyle\sum_{t=1}^{T}\ell^{1/\gamma}_{t}(x)-\min_{u\in\mathcal{X}}\ell_{t}(u)=L^{\mathcal{X}}_{T}(x),$
where the equality uses the fact that $\inf\ell^{1/\gamma}_{t}=\min_{u\in\mathcal{X}}\ell_{t}(u)$ and the last inequality uses $\ell(x)\leq\ell^{1/\gamma}(x)$ from (16). This completes the proof.
C.6 Hardness of achieving an $\mathcal{O}(D\sqrt{G^{\mathcal{X}}_{T}(x)})$ regret upper bound
We show that an $\mathcal{O}(LD^{2}+D\sqrt{G_{T}^{\mathcal{X}}(x)})$ upper bound is also generally not achievable.
Proposition C.1.
Let $\gamma=\Omega(1)$ be some constant. For any $\tau\in[0,\tfrac{1}{2})$, there exists a loss sequence $\{\ell_{t}\}_{t=1}^{T}$ satisfying A1, A2 such that for any online algorithm and $x\in\mathcal{X}$, $\tfrac{\rho_{T}(x)}{LD^{2}+D\sqrt{G_{T}^{\mathcal{X}}(x)}}=\Omega(T^{\tau})$.
Proof.
Consider the lower bound instance with $\mathcal{X}=[-\delta,\delta]$, $D=2\delta$, and $\ell_{t}(x)=g_{t}x$, where $g_{t}=\sigma_{t}$ is a Rademacher random variable. Each $\ell_{t}$ is linear and thus 0-smooth. Then
$\ell_{t}^{1/\gamma}(x)=\min_{z\in[-\delta,\delta]}\{g_{t}z+\tfrac{\gamma}{2}(z-x)^{2}\},$
and for $x\in\mathcal{X}$, we have
$\|\nabla\ell_{t}^{1/\gamma}(x)\|^{2}=\gamma^{2}|\Pi_{[-\delta,\delta]}[x-\tfrac{1}{\gamma}g_{t}]-x|^{2}\leq 4\gamma^{2}\delta^{2}.$
Hence, $G_{T}^{\mathcal{X}}(x)\leq 4\gamma^{2}\delta^{2}T$. The lower bound argument gives $\rho_{T}(x)\geq\Omega(\delta\sqrt{T})$, while the desired $\mathcal{O}(LD^{2}+D\sqrt{G_{T}^{\mathcal{X}}(x)})$ upper bound with $L=0$ yields
$\rho_{T}(x)=\mathcal{O}(\delta\gamma\sqrt{4\delta^{2}T})=\gamma\delta^{2}\sqrt{T}.$
Taking $\delta=T^{-\beta}$, we have lower bound of $\Omega(T^{\frac{1}{2}-\beta})$ and an upper bound of $\mathcal{O}(\gamma T^{\frac{1}{2}-2\beta})$ that contradict each other.